### v1

# Black-box Adversaries from Latent Space: Unnoticeable Attacks on Human Pose and Shape Estimation

Abstract
Expressive human pose and shape (EHPS) estimation is vital for digital human generation, particularly in live-streaming applications. However, most existing EHPS models focus primarily on minimizing estimation errors, with limited attention on potential security vulnerabilities. Current adversarial attacks on EHPS models often require white-box access (e.g., model details or gradients) or generate visually conspicuous perturbations, limiting their practicality and ability to expose real-world security threats. To address these limitations, we propose a novel Unnoticeable Black-Box Attack (UBA) against EHPS models. UBA leverages the latent-space representations of natural images to generate an optimal adversarial noise pattern and iteratively refine its attack potency along an optimized direction in digital space. Crucially, this process relies solely on querying the model’s output, requiring no internal knowledge of the EHPS architecture, while guiding the noise optimization toward greater stealth and effectiveness. Extensive experiments and visual analyses demonstrate the superiority of UBA. Notably, UBA increases the pose estimation errors of EHPS models by 17.27%–58.21% on average, revealing critical vulnerabilities. These findings underscore the urgent need to address and mitigate security risks associated with digital human generation systems.
1 Introduction
Expressive human pose and shape (EHPS) estimation from monocular images or videos is fundamental to digital human modeling and underpins key applications such as live streaming, virtual reality (VR), and interactive gaming [26, 31, 46, 42, 2, 13, 12]. In these scenarios, EHPS systems play a critical role in animating realistic digital avatars capable of dynamically replicating human movements and facial expressions in real time. Although recent advancements have significantly enhanced estimation accuracy, the security vulnerabilities of such models remain largely underexplored. Growing evidence [39, 3, 33, 45] indicate that these systems are susceptible to adversarial perturbations; however, existing attack methods against EHPS [23] often depend on access to internal model parameters or gradient information and typically produce visually noticeable artifacts. These limitations not only hinder the practical deployment of such attacks but also contribute to the limited attention this issue has received within the research community.
Existing EHPS studies primarily focus on reducing errors in body pose estimation [46, 42] and facial pose estimation [18, 28, 24, 2]. This is typically achieved using parametric human body models such as SMPL-X [31] or SMPLer-X [2], which leverage deep neural networks to capture the complex geometry and motion of the human body. While these models exhibit strong performance, their robustness against adversarial attacks has received limited attention. TBA [23] was the first to propose a white-box transferable adversarial attack against EHPS. It utilizes the gradient of the SMPLer-X model to generate adversarial examples that significantly increase the estimation error across various EHPS models, thereby exposing potential security vulnerabilities. However, the perturbations produced by TBA are visually perceptible and fail to pass human inspection, which limits their applicability in real-world scenarios. Despite their effectiveness, such attacks have yet to attract sufficient attention from the research community. Considering these factors, we argue that an ideal adversarial attack on EHPS should meet three key criteria: i) the perturbation must be visually imperceptible; ii) the required prior knowledge of the attacker should be minimal, aligning with realistic threat models; iii) the adversarial examples must achieve a high error growth rate. However, a substantial gap remains between current methods and these ideal standards.
To address this gap, we propose a novel Unnoticeable Black-box Attack (UBA) targeting EHPS models. Inspired by [6, 32], UBA perturbs the latent space of a pretrained variational autoencoder (VAE) [20, 11] by injecting low-magnitude Gaussian noise. Leveraging the nonlinear generative capabilities of the decoder of VAE [19, 37], this will produce an initial adversarial example that is visually indistinguishable from the original input. The resulting pixel-wise difference is treated as an initial perturbation and subsequently refined through an iterative optimization process guided by a customized multi-task loss function. This loss simultaneously maximizes the discrepancy in the EHPS model’s predictions while minimizing perceptual deviation from the clean image. Notably, the attack is conducted under a strict query budget, ensuring both query efficiency and imperceptibility. By exploiting latent-space amplification and aligning perturbations with task-relevant features, UBA achieves high attack efficacy while maintaining visual stealth. In conclusion, our main contributions are:
$\bullet$ We propose the first black-box adversarial attack tailored to EHPS systems, satisfying realistic constraints and highlighting critical vulnerabilities for digital human generation.
$\bullet$ By injecting noise into the VAE latent space, we generate visually imperceptible adversarial examples that exploit the decoder’s nonlinear generative capabilities.
$\bullet$ A custom multi-task loss balances attack strength and imperceptibility, with iterative refinement performed under a strict query limit to avoid detection.
$\bullet$ Extensive experiments demonstrate that our method significantly degrades pose estimation performance, increasing EHPS models’ pose estimation errors by 17.27%–58.21% on average.
2 Related Work
Expressive Human Pose and Shape Estimation. Expressive Human Pose and Shape Estimation (EHPS) is fundamental to digital human modeling, with applications in virtual environments, live broadcasting, avatars, and animation [40, 22, 30, 47, 29, 34]. Recent years have seen significant advances in improving the accuracy and realism of EHPS models. Pavlakos et al. [31] introduced SMPL-X, a comprehensive 3D model incorporating articulated hands and expressive facial features for improved monocular pose capture. Moon et al. [28] proposed Hand4Whole, leveraging MCP joint features to predict wrist rotations precisely while excluding full-body context for better finger estimation. Lin et al. [24] developed OSX, a one-stage framework using a Component Aware Transformer to reconstruct whole-body meshes without post-processing. Cai et al. [2] presented SMPLer-X, a generalist foundation model trained on diverse datasets, achieving strong generalization and competitive performance across EHPS benchmarks. However, existing research has predominantly concentrated on enhancing estimation accuracy, with limited attention to the models’ vulnerability under attack threats.
Adversarial Attack against Deep Neural Networks. Adversarial attacks are commonly employed to assess the robustness of deep neural networks (DNNs) against malicious inputs [16, 9, 34, 14, 15, 44]. These attacks introduce subtle perturbations that mislead models while preserving the visual appearance of the input. Huang et al. [17] proposed TT3D, which integrates NeRF-based multi-view reconstruction with dual optimization of texture features and MLP parameters to generate natural-looking, transferable 3D adversarial examples. Chen et al. [5] presented the first content-aware, unrestricted adversarial attack based on diffusion models, enhancing transferability and realism by optimizing semantic perturbations in latent space. Chen et al. [4] proposed DiffAttack, a diffusion-based framework that achieves highly transferable and imperceptible black-box attacks by jointly modulating cross-attention and self-attention mechanisms. Despite these advancements, the robustness of EHPS models remains largely underexplored. Li et al. [23] proposed TBA consisting of a dual heterogeneous noise generator, the first EHPS-targeted attack; however, its visible perturbations and reliance on gradient access limit its practicality. Consequently, TBA has failed to draw sufficient attention from the research community regarding EHPS security risks. To address this gap, we propose a unnoticeable black-box attack that reveals critical vulnerabilities in EHPS models.
3 Preliminaries
Threat Model. This paper investigates black-box adversarial attack techniques targeting EHPS models. Specifically, the attacker lacks access to any details of the model’s training process, including the dataset, architecture, and parameters, and is only permitted a limited number of queries via the model’s API. The generated adversarial examples are directly submitted to the model for inference. We aim to raise awareness in the research community about security risks in the domain of digital humans. Other attack types with limited practical applicability (e.g., training-controlled or model modification attacks) are excluded from consideration, as they require additional attacker capabilities and are thus unsuitable for this study.
The Main Pipeline of EHPS. Expressive Human Pose and Shape (EHPS) estimation aims to reconstruct detailed 3D representations of the human body, face, and hands from monocular images. This capability underpins a wide array of applications, including virtual environments, live broadcasting, avatar generation, and animation. Formally, the ground-truth pose is denoted as $\alpha\in\Phi^{53\times 3}$, comprising the body pose $\alpha_{\text{body}}\in\Phi^{22\times 3}$, left- and right-hand poses $\alpha_{\text{lhand}},\alpha_{\text{rhand}}\in\Phi^{15\times 3}$, and the jaw pose $\alpha_{\text{jaw}}\in\Phi^{1\times 3}$. Additionally, individual-specific geometry and facial expressions are encoded by the shape vector $\beta\in\Phi^{10}$ and the expression vector $\gamma\in\Phi^{10}$, respectively. Let $\mathcal{P}$ denote the EHPS model; given a unseen monocular image $x^{{}^{\prime}}$, the objective is to estimate the complete set of pose, shape, and expression parameters $\left\{\alpha^{{}^{\prime}},\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right\}$, e.g., $\mathcal{P}(x^{{}^{\prime}})=\left\{\alpha^{{}^{\prime}},\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right\}$. The estimation process minimizes the parameter reconstruction loss:
$\min_{{\alpha^{{}^{\prime}},\beta^{{}^{\prime}},\gamma^{{}^{\prime}}}}\left(\|\alpha^{{}^{\prime}}-\alpha\|+\|\beta^{{}^{\prime}}-\beta\|+\|\gamma^{{}^{\prime}}-\gamma\|\right).$ (1)
Finally, 3D joints are rendered using the SMPL-X function $\mathcal{S}$ and a joint regressor $\mathcal{R}$, denoted as $\mathcal{R}(\mathcal{S}(\alpha^{{}^{\prime}},\beta^{{}^{\prime}},\gamma^{{}^{\prime}}))$.
4 Unnoticeable Black-box Attack (UBA)
4.1 Problem Formulation
As noted in previous sections, existing EHPS models primarily aim to reduce human pose and shape estimation errors, with limited emphasis on their vulnerability to attacks. Although prior work [23] introduced the first adversarial attack against EHPS, exposing its inherent fragility, its real-world applicability remains constrained due to the visible nature of adversarial perturbations and the unrealistic assumption that attackers have access to internal model information. These limitations hinder the community’s full recognition of the intrinsic risks associated with EHPS. To address this, we propose an unnoticeable black-box attack tailored to a realistic setting, where the attacker can only interact with the model through query outputs. To avoid detection by the owners of the EHPS model due to excessive queries, the attack is carried out under a strict query budget.
Attacker’s Goals. The UBA is designed with two primary objectives: 1) effectiveness and 2) unnoticeability. Specifically, the effectiveness demands that the adversarial examples significantly increase errors in human pose and shape estimation. The unnoticeability ensures that EHPS model owners are unable to identify the adversarial examples. The formalization is as follows:
$\mathcal{P}(\hat{x})=\left\{\hat{\alpha}\in\Phi^{53\times 3},\hat{\beta}\in\Phi^{10},\hat{\gamma}\in\Phi^{10}\right\},$ (2)
$\max_{{\hat{\alpha},\hat{\beta},\hat{\gamma}}}\left(\|\hat{\alpha}-\alpha\|+\|\hat{\beta}-\beta\|+\|\hat{\gamma}-\gamma\|\right),\quad\text{s.t.}\quad\left\|x-\hat{x}\right\|\leq\varepsilon,$ (3)
where $x$ denotes a clean image, $x^{adv}$ represents the corresponding adversarial sample, and $\varepsilon$ indicates the perturbation threshold. $\left\{\alpha,\beta,\gamma\right\}$ denotes the ground-truth human pose and shape parameters.
Figure 1: Overview of the Unnoticeable Black-box Attack (UBA) pipeline. The framework consists of two stages: a) noise injection in latent space, and b) noise enhancement.
4.2 The proposed Method
As shown in Fig. 1, our proposed UBA method is a black-box attack framework designed for EHPS models, consisting of two primary stages: 1) injecting random noise into the latent space; and 2) optimizing the noise to maximize attack effectiveness while remaining unnoticeable. Notably, our goal is to reveal vulnerabilities in EHPS models and raise awareness of their security risks, not to cause harm.
Noise Injection in Latent Space. Motivated by the insights of Shukla et al. [35], who challenged the necessity of norm-bounded constraints in pixel space, we argue that adversarial perturbations are more effective when applied in the latent space. Unlike conventional methods that manipulate input images directly [8, 27, 5, 4], this strategy targets internal representations that more directly govern model predictions. Importantly, perturbing the latent space allows for adversarial behavior while preserving high perceptual fidelity in the image space. To reduce attack-side burden and increase real-world applicability, we directly utilize a publicly available pretrained generative model, without any additional training or fine-tuning. To this end, we employ an efficient, pretrained Variational Autoencoder (VAE) [20] to generate adversarial samples.
The VAE is trained by maximizing the evidence lower bound and utilizes the reparameterization trick to enable differentiable sampling:
$\mathcal{L}(\phi,\theta;x)=\mathbb{E}_{q_{\phi}(z|x)}[\log p_{\theta}(x|z)]-D_{\mathrm{KL}}(q_{\phi}(z|x),p(z)),$ (4)
$z=\mu_{\phi}(x)+\sigma_{\phi}(x)\odot\epsilon_{1},\quad\epsilon_{1}\sim\mathcal{N}(0,I),$ (5)
where $\phi$ and $\theta$ denote the encoder and decoder parameters, respectively. $p(z)=\mathcal{N}(0,I)$ is the standard Gaussian prior. $\mathcal{G}_{\phi}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n},z\mapsto x$, $d$ and $n$ denote the dimension of the latent space and the pixel space. This design yields a smooth and continuous latent space [21].
Building on this property, we inject an additional Gaussian noise $\epsilon_{2}$ into the latent representation $z$, creating a variationally grounded sampling-based perturbation that preserves the Gaussianity of the distribution. Specifically, we define $\hat{z}=z+\eta\cdot\epsilon_{2}$, where $\eta$ is a hyperparameter controlling the perturbation magnitude, so that $p(\hat{z})\sim\mathcal{N}(0,I)$. When decoded by the Decoder $\mathcal{G}_{\theta}$, the perturbed latent representation $\hat{z}$ yields outputs $\hat{x}=\mathcal{G}_{\theta}(\hat{z})$. We perform a first-order Taylor expansion of $\mathcal{G}_{\theta}(z)$:
$\mathcal{G}_{\theta}(z+\eta\cdot\epsilon_{2})=\mathcal{G}_{\theta}(z)+J_{\mathcal{G}_{\theta}}(z)\cdot(\eta\cdot\epsilon_{2})+o(||\eta\cdot\epsilon_{2}||),$ (6)
where $J_{\mathcal{G}_{\theta}}(z)$ denotes the Jacobian of the decoder $\mathcal{G}_{\theta}$ at $z$. By omitting the higher-order terms $o(||\eta\cdot\epsilon_{2}||)$, we can obtain:
$\hat{x}\approx\mathcal{G}_{\theta}(z)+J_{\mathcal{G}_{\theta}}(z)\cdot(\eta\cdot\epsilon_{2}).$ (7)
Consequently, the perturbation $\Delta x=\hat{x}-x$ in the pixel space can be expressed as:
$\Delta x\approx J_{\mathcal{G}_{\theta}}(z)\cdot(\eta\cdot\epsilon_{2}).$ (8)
Specifically, the perturbation $\Delta x$ in the pixel space exhibits a linear relationship with the perturbation $\eta\cdot\epsilon_{2}$ in the latent space. Since the pre-trained model $\mathcal{G}_{\theta}$ is typically continuously differentiable, the magnitude of the perturbation is consequently constrained as follows:
$\Delta x\approx J_{\mathcal{G}_{\theta}}(z)\cdot(\eta\cdot\epsilon_{2})\leq\eta\cdot||J_{\mathcal{G}_{\theta}}(z)||_{2}\cdot||\epsilon_{2}||.$ (9)
If the perturbation magnitude $\eta$ is sufficiently small, the corresponding distortion in pixel space remains imperceptible, even in directions associated with large Jacobian values.
Similarly, a first-order approximation is applied to the target model:
$\mathcal{P}(\hat{x})=\mathcal{P}(x+J_{\mathcal{G}_{\theta}}(z)\cdot(\eta\cdot\epsilon_{2}))\approx\mathcal{P}(x)+J_{\mathcal{P}}(x)\cdot J_{\mathcal{G}_{\theta}}(z)\cdot(\eta\cdot\epsilon_{2}).$ (10)
We define the threshold for significant changes in the estimation parameters of the EHPS model as $\delta$. If the deviation exceeds this threshold, the model is considered to be misled, i.e., $||\mathcal{P}(\hat{x})-\mathcal{P}(x)||\geq\delta$. According to Eq. 10, this leads to the inequality $J_{\mathcal{P}}(x)\cdot J_{\mathcal{G}_{\theta}}(z)\cdot(\eta\cdot\epsilon_{2})\geq\delta$.
By combining the Jacobian matrices into a single term $\mathbb{J}=J_{\mathcal{P}}(x)\cdot J_{\mathcal{G}_{\theta}}(z)$, the inequality can be reformulated as $||\mathbb{J}\cdot(\eta\cdot\epsilon_{2})||\geq\delta$. Given that $\epsilon_{2}\sim\mathcal{N}(0,I)$, the expected perturbation is:
$\mathbb{E}[||\mathbb{J}\cdot(\eta\cdot\epsilon_{2})||^{2}_{2}]=\eta^{2}\mathbb{E}[\epsilon_{2}^{\top}\mathbb{J}^{\top}\mathbb{J}\epsilon_{2}]=\eta^{2}\text{tr}(\mathbb{J}^{\top}\mathbb{J}).$ (11)
To minimize the perturbation magnitude $\eta$ while effectively crossing the decision boundary, the perturbation should be aligned with the right singular vector corresponding to the largest singular value:
$\mathbb{J}=U\Sigma V^{\top},\quad\Sigma=\text{diag}(\sigma_{1},\sigma_{2},\ldots),\quad\sigma_{1}\geq\sigma_{2}\geq\cdots$ (12)
The most sensitive direction corresponds to the first column vector $v_{1}$ of $V$. If $\epsilon_{2}$ is set to $\upsilon_{1}$, the condition for the minimum perturbation magnitude becomes:
$||\mathbb{J}\cdot(\eta\cdot\upsilon_{1})||_{2}=\eta\cdot\sigma_{1}\geq\delta\Rightarrow\eta\geq\frac{\delta}{\sigma_{1}}.$ (13)
In summary, an attacker can amplify variations in sensitive features by injecting small perturbations into the latent space, causing the prediction to deviate rapidly from its original stable state and leading to significant errors.
Noise Enhancement. Inspired by Projected Gradient Descent (PGD) [27], we adopt an iterative refinement strategy to improve the effectiveness of adversarial examples, guided by feedback from the EHPS model. However, updating the noise in the latent space entails significant computational overhead. Specifically, the latent perturbation is defined as:
$\hat{z}^{(t+1)}=\hat{z}^{(t)}+\xi_{z}\nabla_{\hat{z}^{(t)}}\mathcal{L}(\mathcal{P}(\mathcal{G}_{\theta}(\hat{z}^{(t)})),(\alpha,\beta,\gamma)),$ (14)
$\nabla_{\hat{z}^{(t)}}\mathcal{L}(\mathcal{P}(\mathcal{G}_{\theta}(\hat{z}^{(t)})),(\alpha,\beta,\gamma))=\frac{\partial\mathcal{L}}{\partial\mathcal{P}}\frac{\partial\mathcal{P}}{\partial\mathcal{G}_{\theta}}\frac{\partial\mathcal{G}_{\theta}}{\partial\hat{z}^{(t)}},$ (15)
where $t$ denotes the iteration steps, $\hat{z}^{(t)}$ represents the latent perturbation at step $t$, and $\xi_{z}$ is the learning rate for the gradient update in latent space. The term $\frac{\partial\mathcal{G}_{\theta}}{\partial\hat{z}^{(t)}}$ denotes the Jacobian matrix of the Decoder, whose computation is prohibitively expensive due to its high dimensionality ($n\times d$, often very large). Furthermore, each update necessitates a complete forward pass through the decoder network, resulting in substantial memory and computational consumption. Consequently, performing iterative updates directly in the latent space is highly inefficient (see Method C in Table 3).
To address the aforementioned issue, we propose enhancing the noise directly in the pixel space. The initial pixel-space perturbation is defined as:
$\Delta x^{(0)}=x^{(0)}-x,\quad\Delta x^{(0)}=\mathcal{G}_{\theta}(z+\eta\cdot\epsilon_{2}).$ (16)
We perform iterative updates to the perturbation in pixel space as follows:
$\Delta x^{(t+1)}=\Delta x^{(t)}+\xi_{x}\cdot\nabla_{\Delta x^{(t)}}\mathcal{L}(\mathcal{P}(x+\Delta x^{(t)}),(\alpha,\beta,\gamma)),$ (17)
$\nabla_{\Delta x^{(t)}}\mathcal{L}(\mathcal{P}(x+\Delta x^{(t)}),(\alpha,\beta,\gamma))=\frac{\partial\mathcal{L}}{\partial\mathcal{P}}\frac{\partial\mathcal{P}}{\partial(x+\Delta x^{(t)})},$ (18)
where $\xi_{x}$ denotes the learning rate for gradient updates in pixel space. Importantly, computing gradients in pixel space does not require backpropagation through the decoder network. Instead, perturbations are applied directly to the original image $x$, thereby eliminating the need to compute the high-dimensional Jacobian transformation from latent to pixel space. This approach substantially reduces both the computational complexity and memory usage of each gradient update step.
Furthermore, to ensure strong adversarial strength while maintaining the perturbation’s imperceptibility, we incorporate a regularization term into the optimization process. To this end, we introduce a novel multi-task loss function to jointly achieve these objectives:
$\mathcal{L}=-\mathbb{E}_{(\hat{\alpha},\alpha,\hat{\beta},\beta,\hat{\gamma},\gamma)}\left[\|\hat{\alpha}-\alpha\|^{2}+\|\hat{\beta}-\beta\|^{2}+\|\hat{\gamma}-\gamma\|^{2}\right]+\frac{1}{w\times h}\sum_{i=0}^{w-1}\sum_{j=0}^{h-1}\left\|\hat{x}_{i,j}-x_{i,j}\right\|^{2},$ (19)
where $w,h$ denotes the width and height of $x$. The first term enforces the perturbed pose parameters to deviate from the ground-truth parameters (enhancing the attack effectiveness), while the second term strictly constrains pixel-level variations to maintain the imperceptibility of the perturbation. According the Eqs 17, 18 and 19, the update rule for the pixel-space perturbation is formulated as:
$\Delta x^{(t+1)}=\Delta x^{(t)}+\xi_{x}\cdot(\frac{\partial\mathcal{L}}{\partial\mathcal{P}}\frac{\partial\mathcal{P}}{\partial(x+\Delta x^{(t)})}-\lambda\cdot(x^{(t)}-x)),$ (20)
where $\lambda$ is a hyperparameter that balances attack strength and imperceptibility. It is evident that the computational complexity of updates in latent space is $\mathcal{O}(n\times d)$, whereas the complexity of pixel-space updates is only $\mathcal{O}(n)$. Therefore, the proposed noise enhancement strategy in pixel space incurs significantly lower computational overhead (see the Method C and UBA (Ours) in Table 3).
Notably, we interact with the model solely via its application interface (API), without access to internal parameters, and impose a strict limit on the number of queries $t$. This design reduces the risk of detection due to excessive querying and alleviates the burden on the attacker.
5 Experiments
5.1 Experimental Settings
Datasets and EHPS Models. We evaluate our method on two benchmarks: 3DPW [38], with in-the-wild RGB videos and 3D SMPL pose annotations, and UBody [24], a large-scale upper-body mesh dataset with 1.05 million annotated images across 15 scenarios. For black-box evaluation, we test UBA on state-of-the-art EHPS models, including Hand4Whole [28], OSX [24], and SMPLer-X [2], which provides four ViT-based pre-trained variants, denoted as SMPLer-X-M, where M denotes the size of the Vision Transformer (ViT) backbone [1] (M $\in$ S, B, L, H).
Baseline and Evaluation Metrics. Currently, only one adversarial attack targets EHPS. To evaluate UBA, we compare it with state-of-the-art methods in computer vision, including classical attacks (FGSM [8], PGD [27]), diffusion-based approaches (ACA [5], DiffAttack [4]), and the EHPS-specific TBA [23]. We report MPJPE and MPVPE for 3D joint and mesh accuracy, and PA-MPJPE/PA-MPVPE for alignment-aware evaluation, following prior EHPS benchmarks [28, 24, 2]. All metrics are in millimeters (mm).
Implementation Details. The experiments are conducted on an Ubuntu 20.04.6 LTS system using the PyTorch framework, a single NVIDIA L20 GPU, and 45 GB of RAM. The VAE model employed is stabilityai/sd-vae-ft-mse from Hugging Face  [^1] An open source platform. https://huggingface.co/stabilityai/sd-vae-ft-mse. We set the number of queries to $t=3$, the perturbation magnitude to $\eta=0.05$, and the balancing hyperparameter to $\lambda=0.5$.
5.2 Attack Performance on EHPS Models
Table 1: Performance comparison of different adversarial attack methods on state-of-the-art EHPS models on the UBody dataset. The error growth rates are marked in gray. The maximum error and maximum error growth rate on each setting are highlighted underlined.
Model Attack PA MPVPE $\downarrow$ (mm) MPVPE $\downarrow$ (mm) PA MPJPE $\downarrow$ (mm)
All Hands Face All Hands Face Body Hands
SMPLer-X-H Clean 24.64 8.47 2.24 41.22 30.88 16.62 29.29 8.64
FGSM 29.83 (21.06%) 9.03 (6.61%) 2.41 (7.59%) 57.76 (40.13%) 36.03 (16.68%) 20.11 (21.00%) 34.82 (18.88%) 9.18 (6.25%)
PGD 29.59 (20.09%) 9.17 (8.26%) 2.49 (11.16%) 60.06 (45.71%) 34.50 (11.72%) 22.00 (32.37%) 34.55 (17.96%) 9.33 (7.99%)
ACA 26.84 (8.93%) 9.14 (7.91%) 2.40 (7.14%) 46.29 (12.30%) 34.12 (10.49%) 18.04 (8.54%) 31.39 (7.17%) 9.32 (7.87%)
DiffAttack 28.54 (15.83%) 9.32 (10.04%) 2.43 (8.48%) 53.41 (29.57%) 35.41 (14.67%) 19.70 (18.53%) 33.80 (15.40%) 9.48 (9.72%)
TBA 30.00 (21.75%) 9.09 (7.32%) 2.43 (8.48%) 58.13 (41.02%) 36.29 (17.52%) 21.19 (27.50%) 35.36 (20.72%) 9.25 (7.06%)
UBA (Ours) 43.94 (78.33%) 9.90 (16.88%) 2.80 (25.00%) 88.60 (114.94%) 48.69 (57.69%) 28.38 (70.76%) 50.48 (72.35%) 10.08 (16.67%)
SMPLer-X-L Clean 25.66 8.98 2.39 43.26 33.11 17.54 30.14 9.16
FGSM 32.88 (28.14%) 9.32 (3.79%) 2.51 (5.02%) 63.99 (47.92%) 39.44 (19.12%) 22.52 (28.39%) 38.58 (28.00%) 9.49 (3.60%)
PGD 30.82 (20.11%) 9.36 (4.23%) 2.62 (9.62%) 63.07 (45.79%) 36.61 (10.57%) 22.84 (30.22%) 36.40 (20.77%) 9.52 (3.93%)
ACA 28.13 (9.63%) 9.69 (7.91%) 2.52 (5.44%) 49.39 (14.17%) 36.47 (10.15%) 19.18 (9.35%) 32.50 (7.83%) 9.86 (7.64%)
DiffAttack 30.65 (19.45%) 9.74 (8.46%) 2.49 (4.18%) 57.57 (33.08%) 37.96 (14.65%) 20.53 (17.05%) 36.32 (20.50%) 9.90 (8.08%)
TBA 33.50 (30.55%) 9.43 (5.01%) 2.69 (12.55%) 66.16 (52.94%) 40.50 (22.32%) 22.05 (25.71%) 39.01 (29.73%) 9.60 (4.80%)
UBA (Ours) 47.86 (86.52%) 9.83 (9.47%) 2.92 (22.18%) 91.46 (114.42%) 52.17 (57.57%) 33.01 (88.20%) 54.50 (80.82%) 10.03 (9.50%)
SMPLer-X-B Clean 28.95 9.72 2.60 50.75 38.02 19.82 33.43 9.91
FGSM 38.06 (31.47%) 10.06 (3.50%) 2.83 (8.85%) 78.25 (54.19%) 45.42 (19.46%) 25.47 (28.51%) 43.83 (31.11%) 10.24 (3.33%)
PGD 34.56 (19.38%) 9.85 (1.34%) 2.87 (10.38%) 71.73 (41.34%) 41.46 (9.05%) 24.47 (23.46%) 40.16 (20.13%) 10.03 (1.21%)
ACA 31.48 (8.74%) 10.11 (4.01%) 2.74 (5.38%) 57.02 (12.35%) 41.31 (8.65%) 21.99 (10.95%) 36.16 (8.17%) 10.30 (3.94%)
DiffAttack 32.04 (10.67%) 10.12 (4.12%) 2.72 (4.62%) 58.23 (14.74%) 41.48 (9.10%) 21.79 (9.94%) 37.18 (11.22%) 10.30 (3.94%)
TBA 37.44 (29.33%) 9.80 (0.82%) 2.90 (11.54%) 76.41 (50.56%) 46.25 (21.65%) 25.63 (29.31%) 42.69 (27.70%) 10.04 (1.31%)
UBA (Ours) 53.38 (84.39%) 10.61 (9.16%) 3.23 (24.23%) 106.97 (110.78%) 57.10 (50.18%) 36.21 (82.69%) 62.12 (85.82%) 10.82 (9.18%)
SMPLer-X-S Clean 32.28 10.05 2.83 57.23 42.80 22.26 37.38 10.27
FGSM 40.02 (23.98%) 10.73 (6.77%) 3.01 (6.36%) 82.44 (44.05%) 49.33 (15.26%) 29.30 (31.63%) 46.44 (24.24%) 10.96 (6.72%)
PGD 36.90 (14.31%) 10.63 (5.77%) 3.07 (8.48%) 78.03 (36.43%) 45.79 (6.99%) 26.85 (20.62%) 43.37 (16.02%) 10.89 (6.04%)
ACA 34.48 (6.82%) 10.53 (4.78%) 2.93 (3.53%) 61.99 (8.32%) 45.04 (5.23%) 23.63 (6.15%) 39.66 (6.10%) 10.72 (4.38%)
DiffAttack 34.27 (6.16%) 10.54 (4.88%) 2.89 (2.12%) 61.31 (7.13%) 45.25 (5.72%) 23.47 (5.44%) 39.43 (5.48%) 10.74 (4.58%)
TBA 39.52 (22.43%) 10.96 (9.05%) 3.17 (12.01%) 81.48 (42.37%) 49.46 (15.56%) 28.72 (29.02%) 46.06 (23.22%) 11.25 (9.54%)
UBA (Ours) 50.16 (55.39%) 10.53 (4.78%) 3.33 (17.67%) 98.24 (71.66%) 59.06 (37.99%) 34.70 (55.88%) 58.62 (56.82%) 10.78 (4.97%)
OSX Clean 40.85 9.37 3.54 86.36 42.72 25.72 49.26 9.62
FGSM 45.68 (11.82%) 9.86 (5.23%) 3.29 (-7.06%) 104.51 (21.02%) 46.71 (9.34%) 33.64 (30.84%) 54.17 (9.97%) 10.13 (5.30%)
PGD 42.47 (3.97%) 9.68 (3.31%) 3.00 (-15.25%) 97.86 (13.32%) 43.89 (2.74%) 28.78 (11.94%) 51.11 (3.76%) 9.95 (3.43%)
ACA 41.93 (2.64%) 9.66 (3.09%) 3.48 (-1.69%) 89.10 (3.17%) 45.02 (5.38%) 27.05 (5.21%) 49.92 (1.34%) 9.94 (3.33%)
DiffAttack 40.66 (-0.47%) 9.67 (3.20%) 3.11 (-12.15%) 88.86 (2.89%) 43.99 (2.97%) 26.04 (1.28%) 48.64 (-1.26%) 9.95 (3.43%)
TBA 41.37 (1.27%) 9.67 (3.20%) 2.98 (-15.82%) 97.25 (12.61%) 45.37 (6.20%) 28.05 (9.10%) 49.57 (0.63%) 9.94 (3.33%)
UBA (Ours) 60.36 (47.76%) 10.44 (11.42%) 3.86 (9.04%) 134.39 (55.62%) 55.82 (30.66%) 47.73 (85.65%) 75.17 (52.60%) 10.73 (11.54%)
Hand4Whole Clean 41.04 7.85 2.96 93.81 35.60 30.30 49.80 7.97
FGSM 41.03 (-0.02%) 7.85 (0.00%) 2.96 (0.00%) 93.82 (0.01%) 35.61 (0.03%) 30.30 (0.00%) 49.80 (0.00%) 7.97 (0.00%)
PGD 41.06 (0.05%) 7.85 (0.00%) 2.96 (0.00%) 93.87 (0.06%) 35.62 (0.06%) 30.31 (0.03%) 49.82 (0.04%) 7.97 (0.00%)
ACA 41.40 (0.88%) 7.83 (-0.25%) 2.98 (0.68%) 95.15 (1.43%) 36.02 (1.18%) 30.71 (1.35%) 50.34 (1.08%) 7.95 (-0.25%)
DiffAttack 41.10 (0.15%) 7.99(1.78%) 2.91 (-1.69%) 94.79 (1.04%) 36.02 (1.18%) 35.11 (15.87%) 50.12 (0.64%) 8.10 (1.63%)
TBA 43.72 (6.53%) 8.10 (3.18%) 3.00 (1.35%) 104.31 (11.19%) 36.57 (2.72%) 34.52 (13.93%) 53.22 (6.87%) 8.22 (3.14%)
UBA (Ours) 56.59 (37.89%) 9.74 (24.08%) 3.52 (18.92%) 128.57 (37.05%) 54.02 (51.74%) 41.83 (38.05%) 71.67 (43.92%) 9.93 (24.59%)
Figure 2: Visualizing various adversarial samples for digital human generation.
We first evaluate the effectiveness of UBA on the UBody dataset, which includes full-body, hand, and facial regions of digital humans. As shown in Table 1, across all tasks, the adversarial examples generated by UBA cause the most significant degradation in the accuracy of all EHPS models, compared to other attack methods. Specifically, the average increase in estimation error ranged from 25.13% to 58.21%. Notably, for the SMPLer-X-H model, the MPVPE (All) increased from 41.22 mm to 88.60 mm, corresponding to a 114.94% error growth rate.
Table 2: Performance comparison of different adversarial attack methods on state-of-the-art EHPS models on the 3DPW dataset. The error growth rates are marked in gray. The maximum error and maximum error growth rate on each setting are highlighted underlined.
Model Attack MPJPE (Body) $\downarrow$ (mm) PA MPJPE (Body) $\downarrow$ (mm)
SMPLer-X-H Clean 75.01 50.57
FGSM 84.40 (12.52%) 56.04 (10.82%)
PGD 88.70 (18.25%) 57.17 (13.05%)
ACA 77.30 (3.05%) 52.47 (3.76%)
DiffAttack 80.46 (7.27%) 54.26 (7.30%)
TBA 88.43 (17.89%) 59.39 (17.44%)
UBA (Ours) 103.45 (37.91%) 65.40 (29.33%)
SMPLer-X-L Clean 75.85 50.67
FGSM 85.05 (12.13%) 56.20 (10.91%)
PGD 90.16 (18.86%) 57.30 (12.12%)
ACA 78.53 (3.53%) 53.02 (4.64%)
DiffAttack 78.70 (3.76%) 53.25 (5.09%)
TBA 91.63 (20.80%) 60.54 (19.48%)
UBA (Ours) 104.53 (37.81%) 64.94 (28.16%)
SMPLer-X-B Clean 79.46 52.62
FGSM 94.04 (18.35%) 60.97 (15.87%)
PGD 94.45 (18.86%) 59.00 (12.12%)
ACA 82.68 (4.05%) 55.60 (5.66%)
DiffAttack 82.93 (4.37%) 56.31 (7.01%)
TBA 97.50 (22.70%) 63.55 (20.77%)
UBA (Ours) 117.97 (48.46%) 73.94 (40.52%)
SMPLer-X-S Clean 82.67 56.65
FGSM 100.05 (21.02%) 65.32 (15.30%)
PGD 98.65 (19.33%) 63.19 (11.54%)
ACA 85.52 (3.45%) 58.86 (3.90%)
DiffAttack 85.61 (3.56%) 58.97 (4.10%)
TBA 108.15 (30.82%) 69.63 (22.91%)
UBA (Ours) 117.78 (42.47%) 74.74 (31.93%)
OSX Clean 94.32 63.87
FGSM 102.09 (8.24%) 68.13 (6.67%)
PGD 99.58 (5.58%) 66.01 (3.35%)
ACA 98.05 (3.95%) 66.39 (3.95%)
DiffAttack 97.78 (3.67%) 66.10 (3.49%)
TBA 97.33 (3.19%) 65.32 (3.49%)
UBA (Ours) 116.89 (23.93%) 75.34 (17.96%)
Hand4Whole Clean 100.66 67.93
FGSM 100.66 (0.00%) 67.93 (0.00%)
PGD 100.66 (0.00%) 67.93 (0.00%)
ACA 100.70 (0.04%) 67.95 (0.03%)
DiffAttack 100.54 (-0.12%) 67.68 (-0.37%)
TBA 115.07 (14.32%) 74.18 (9.20%)
UBA (Ours) 128.72 (27.88%) 80.10 (17.92%)
(a) UBody
(b) PW3D
Figure 3: Attack performance of different perturbation magnitude $\eta$.
(a) UBody
(b) PW3D
Figure 4: Attack performance of different query times $t$.
provided in the Appendix B). As shown in Fig. 4, the number of queries $t$ is positively correlated with the mean error growth rate, suggesting that increasing the number of queries enhances attack effectiveness. Notably, when $t=9$, the mean error growth rates for SMPLer-X-B, SMPLer-X-L, and SMPLer-X-H all exceed 100%. However, an excessive number of queries raises the attacker’s
We further evaluate UBA on the 3DPW dataset, focusing on how adversarial examples impact the estimation of body parts across different EHPS models. As presented in Table 2, UBA consistently causes the greatest reduction in model accuracy. In particular, UBA yielded a minimum error growth rate of 17.92% (increasing PA MPJPE from 67.93 mm to 80.10 mm on the Hand4Whole model) and a maximum of 48.46% (raising MPJPE from 79.46 mm to 118.00 mm on the SMPLer-X-B model), with average increases ranging from 17.27% to 44.49%.
These results demonstrate that UBA is highly effective and destructive. Remarkably, it functions as a black-box attack, requiring minimal prior knowledge—achieving success through API queries alone. Interestingly, we observe negative error growth rates in a few instances, such as PA MPVPE (Face) under the OSX model on UBody, and several other metrics under specific settings, suggesting that adversarial attacks may not only impair EHPS performance but could, in rare cases, enhance estimation accuracy.
To further illustrate the effectiveness of UBA, we conduct qualitative evaluations by visualizing the digital humans generated from both clean and adversarial samples produced by various attack methods. As shown in Fig. 2, UBA leads to noticeable degradation in pose and shape estimation accuracy, resulting in significant deviations in the synthesized digital humans. These visual results clearly highlight the adverse impact of UBA on the EHPS models’ ability to accurately reconstruct human poses (Please refer to Appendix B for more results).
Table 3: Ablation study of different adversarial attack frameworks on state-of-the-art EHPS models on the UBody dataset. The error growth rates are marked in gray. The maximum error and maximum error growth rate on each setting are highlighted underlined.
Model Method PA MPVPE $\downarrow$ (mm) MPVPE $\downarrow$ (mm) PA MPJPE $\downarrow$ (mm) Memory
All Hands Face All Hands Face Body Hands
SMPLer-X-H Clean 24.64 8.47 2.24 41.22 30.88 16.62 29.29 8.64 -
A 32.41 (31.53%) 8.98 (6.02%) 2.41 (7.59%) 63.43 (53.88%) 38.65 (25.16%) 20.16 (21.30%) 38.29 (30.73%) 9.12 (5.56%) 9.93 GB
B 26.27 (6.62%) 8.58 (1.30%) 2.22 (-0.89%) 47.68 (15.67%) 32.28 (4.53%) 18.07 (8.72%) 31.27 (6.76%) 8.74 (1.16%) 7.00 GB
C Out of Memory $\geq$ 45 GB
UBA (Ours) 43.94 (78.33%) 9.90 (16.88%) 2.80 (25.00%) 88.60 (114.94%) 48.69 (57.69%) 28.38 (70.76%) 50.48 (72.35%) 10.08 (16.67%) 12.83 GB
SMPLer-X-L Clean 25.66 8.98 2.39 43.26 33.11 17.54 30.14 9.16 -
A 35.64 (38.89%) 9.18 (2.23%) 2.48 (3.77%) 69.94 (61.97%) 41.61 (25.67%) 23.45 (33.69%) 42.31 (40.38%) 9.35 (2.07%) 6.20 GB
B 27.53 (7.29%) 8.99 (0.11%) 2.42 (1.26%) 50.45 (16.62%) 34.09 (2.96%) 19.09 (8.84%) 32.31 (7.20%) 9.16 (0.00%) 6.77 GB
C 27.24 (6.16%) 8.99 (0.11%) 2.40 (0.42%) 47.74 (10.36%) 34.21 (3.32%) 18.37 (4.73%) 32.05 (6.34%) 9.17 (0.11%) 43.79 GB
UBA (Ours) 47.86 (86.52%) 9.83 (9.47%) 2.92 (22.18%) 91.46 (114.42%) 52.17 (57.57%) 33.01 (88.20%) 54.50 (80.82%) 10.03 (9.50%) 8.25 GB
SMPLer-X-B Clean 28.95 9.72 2.60 50.75 38.02 19.82 33.43 9.91 -
A 37.01 (27.84%) 9.68 (-0.41%) 2.71 (4.23%) 74.09 (45.99%) 44.58 (17.25%) 24.02 (21.19%) 43.11 (28.96%) 9.86 (-0.50%) 3.45 GB
B 31.48 (8.74%) 9.49 (-2.37%) 2.59 (-0.38%) 56.79 (11.90%) 39.01 (2.60%) 21.56 (8.78%) 37.00 (10.68%) 9.67 (-2.42%) 6.24 GB
C 27.24 (-5.91%) 8.99 (-7.51%) 2.40 (-7.69%) 47.74 (-5.93%) 34.21 (-10.02%) 18.37 (-7.32%) 32.05 (-4.13%) 9.17 (-7.47%) 43.63 GB
UBA (Ours) 53.38 (84.39%) 10.61 (9.16%) 3.23 (24.23%) 106.97 (110.78%) 57.10 (50.18%) 36.21 (82.69%) 62.12 (85.82%) 10.82 (9.18%) 8.32 GB
SMPLer-X-S Clean 32.28 10.05 2.83 57.23 42.80 22.26 37.38 10.27 -
A 38.80 (20.20%) 10.38 (3.28%) 2.85 (0.71%) 76.07 (32.92%) 48.19 (12.59%) 25.81 (15.95%) 45.44 (21.56%) 10.60 (3.21%) 2.73 GB
B 35.74 (10.72%) 10.02 (-0.30%) 2.88 (1.77%) 67.46 (17.88%) 44.61 (4.23%) 23.70 (6.47%) 41.41 (10.78%) 10.26 (-0.10%) 7.04 GB
C 33.89 (4.99%) 10.10 (0.50%) 2.85 (0.71%) 60.28 (5.33%) 44.20 (3.27%) 22.74 (2.16%) 39.21 (4.90%) 10.34 (0.68%) 43.33 GB
UBA (Ours) 50.16 (55.39%) 10.53 (4.78%) 3.33 (17.67%) 98.24 (71.66%) 59.06 (37.99%) 34.70 (55.88%) 58.62 (56.82%) 10.78 (4.97%) 7.47 GB
5.3 Ablation Study
Effects of Different Noise Injection Strategies. To further explore the advantages of UBA, we revisit three alternative and more direct noise injection strategies within the UBA framework: A) adding random noise in the pixel space and iteratively updating it by querying the EHPS model; B) injecting random noise into the latent space after VAE encoding, followed by resampling and decoding to generate adversarial examples; and C) injecting random noise into the VAE latent space, then resampling and decoding, with iterative updating it by querying the EHPS model. We evaluate these strategies on the SMPLer-X model using the UBody dataset and report the GPU memory required for generating adversarial examples. As presented in Table 3, our method achieves strong attack performance while requiring only 7.47GB–12.83GB of GPU memory, indicating high practicality. These results demonstrate the effectiveness and efficiency of our approach in attacking EHPS models without the need for extensive computational resources.
Effects of Different Perturbation Magnitude $\eta$. To gain a clearer understanding of the impact of varying perturbation magnitudes $\eta$ on the EHPS model, we conduct experiments using the SMPLer-X model and report the mean error growth rates across all tasks on the UBody and PW3D datasets, respectively (task-specific results are provided in the Appendix B). As illustrated in Fig. 3, different perturbation magnitudes $\eta$ result in comparable attack effectiveness. This phenomenon may be attributed to the model’s sensitivity to latent-space perturbations being concentrated along specific directions rather than being strictly dependent on the perturbation magnitude. As a result, increasing the magnitude does not substantially improve attack performance.
Effects of Different Query Times $t$. Similarly, we conduct an ablation study on the number of queries to the EHPS model API, denoted as $t$, to examine its impact on the performance of adversarial attacks. Due to space limitations, we report only the mean error growth rates (task-specific results are
cost and may alert the platform or API owner. Therefore, although such attacks may not yield optimal performance, conducting them under a strict query budget supports sustained adversarial behavior.
6 Conclusion
This paper introduces a novel Unnoticeable Black-box Attack (UBA) method. By utilizing a pre-trained Variational Autoencoder (VAE), we exploit the structured and continuous nature of its latent space: input samples are first projected into the latent space, where noise is injected, and then decoded to generate perturbations relative to the original inputs, yielding an optimal adversarial pattern. The noise pattern is iteratively refined under a strict query budget, guided by feedback from the EHPS model. To improve the effectiveness of the attack while preserving imperceptibility, we employ a multi-task loss function. Experimental results confirm that UBA achieves high attack success. While UBA is designed to attack EHPS models, its broader aim is to highlight the potential vulnerabilities of EHPS systems and promote awareness within the research community.
Negative Social Impacts. The experimental results validate the proposed method’s strong applicability and attack effectiveness. However, its potential misuse by malicious actors to compromise security-sensitive applications raises serious concerns about the robustness of digital human generation technologies.
Appendix A Technical Details
Datasets. We evaluate our method on two benchmark datasets: 3DPW [38] and UBody [24]. 3DPW comprises in-the-wild RGB video sequences annotated with 3D SMPL poses, capturing diverse human motions in both indoor and outdoor settings using RGB cameras and inertial measurement units (IMUs). UBody is a large-scale dataset for upper-body mesh recovery, serving as a bridge between full-body reconstruction and real-world applications. It contains 1.05 million high-resolution images from 15 real-world scenarios, annotated with 2D whole-body keypoints, person and hand bounding boxes, and SMPL-X parameters.
EHPS Models. To assess the effectiveness of UBA under black-box conditions, we adopt several state-of-the-art EHPS models. In particular, we utilize Hand4Whole [28], OSX [24], and SMPLer-X [2] in our experiments. SMPLer-X provides four pre-trained foundation models, collectively referred to as “SMPLer-X-M”, where M denotes the size of the Vision Transformer (ViT) backbone [1], categorized as S, B, L, H.
Baseline and Evaluation Metrics. Currently, only one adversarial attack has been specifically designed for EHPS. To rigorously evaluate the effectiveness of UBA, we compare it with several state-of-the-art adversarial attack methods in computer vision. Specifically, we assess classical attack algorithms—FGSM [8] and PGD [27]—alongside two diffusion-based approaches, ACA [5] and DiffAttack [4], as well as the first EHPS-specific attack, TBA [23]. For performance evaluation, we report the Mean Per Joint Position Error (MPJPE) and Mean Per-Vertex Position Error (MPVPE), which quantify the accuracy of 3D joint locations and mesh vertex positions, respectively, following prior EHPS benchmarks [28, 24, 2]. Additionally, we include PA-MPJPE and PA-MPVPE, which further align rotation and scale. All metrics are measured in millimeters (mm).
Implementation Details. The experiments are conducted on an Ubuntu 20.04.6 LTS system using the PyTorch framework, a single NVIDIA L20 GPU, and 45 GB of RAM. The VAE model employed is stabilityai/sd-vae-ft-mse from Hugging Face  [^2] An open source platform. https://huggingface.co/stabilityai/sd-vae-ft-mse. The number of queries is set to $t=3$, the perturbation magnitude to $\eta=0.05$, and the hyperparameter balancing attack strength and imperceptibility to $\lambda=0.5$.
Algorithmic Flow. To provide a clearer understanding of the UBA pipeline, we present a detailed the step of generating adversarial samples; see Algorithm 1.
Algorithm 1 Generation of Adversarial Sample
Input: Original Image $x$, Random Gaussian Noise $\epsilon 1$, $\epsilon_{2}$
Output: Adversarial Sample $\hat{x}$
Setting: Encoder of VAE $\mathcal{E}_{\phi}$, Decoder of VAE $\mathcal{G}_{\theta}$, EHPS model $\mathcal{P}$, Mean value $\mu_{\phi}$, Variance $\sigma_{\phi}^{2}$, Number of iterative queries $N$, Step size $\upsilon$, Noise magnitude $\eta$, Learning rate $\zeta$, Latent representation $z$, The ground-truth human pose and shape parameters $\left\{\alpha,\beta,\gamma\right\}$, Multi-task loss function $\mathcal{L}$
1:  Loading parameters $\phi$, $\theta$ of VAE
2:  $\mu_{\phi},\sigma_{\phi}^{2}=\mathcal{E}_{\phi}(x)$
3:  $z=\mu_{\phi}(x)+\sigma_{\phi}(x)\odot\epsilon_{1}$
4:  $\hat{z}=z+\eta\times\epsilon_{2}$
5:  ${\Delta x}=\mathcal{G}_{\theta}(\hat{z})-x$
6:  for $i$ = 1 to $N$ do
7:     $\hat{x}=x+\Delta x$
8:     $\hat{\alpha},\hat{\beta},\hat{\gamma}=P(\hat{x})$
9:     $\mathcal{L}=\mathcal{L}(\hat{\alpha},\hat{\beta},\hat{\gamma},\hat{x},\alpha,\beta,\gamma,x)$
10:     Update
11:     $\Delta x=\Delta x-\zeta\nabla_{\Delta x}\mathcal{L}$
12:  end for
13:  return $\hat{x}=\Delta x+x$
Appendix B Additional Experiments
Table 4: Performance comparison of different perturbation magnitude $\eta$ on state-of-the-art EHPS models on the UBody dataset. The error growth rates are marked in gray. The maximum error and maximum error growth rate on each setting are highlighted underlined.
Model $\eta$ PA MPVPE $\downarrow$ (mm) MPVPE $\downarrow$ (mm) PA MPJPE $\downarrow$ (mm)
All Hands Face All Hands Face Body Hands
SMPLer-X-H Clean 24.64 8.47 2.24 41.22 30.88 16.62 29.29 8.64
0.05 43.94 (78.33%) 9.90 (16.88%) 2.80 (25.00%) 88.60 (114.94%) 48.69 (57.69%) 28.38 (70.76%) 50.48 (72.35%) 10.08 (16.67%)
0.10 44.46 (80.44%) 9.94 (17.36%) 2.79 (24.55%) 89.22 (116.45%) 49.44 (60.10%) 28.03 (68.65%) 51.03 (74.22%) 10.11 (17.01%)
0.15 44.73 (81.53%) 10.02 (18.30%) 2.82 (25.89%) 90.74 (120.14%) 49.62 (60.69%) 28.02 (68.59%) 51.46 (75.69%) 10.21 (18.17%)
0.20 44.51 (80.64%) 10.13 (19.60%) 2.80 (25.00%) 90.76 (120.18%) 49.76 (61.14%) 28.54 (71.72%) 51.38 (75.42%) 10.31 (19.33%)
0.25 45.43 (84.38%) 9.98 (17.83%) 2.81 (25.45%) 92.81 (125.16%) 49.98 (61.85%) 28.20 (69.68%) 52.79 (80.23%) 10.16 (17.59%)
SMPLer-X-L Clean 25.66 8.98 2.39 43.26 33.11 17.54 30.14 9.16
0.05 47.86 (86.52%) 9.83 (9.47%) 2.92 (22.18%) 91.46 (114.42%) 52.17 (57.57%) 33.01 (88.20%) 54.50 (80.82%) 10.03 (9.50%)
0.10 47.91 (86.71%) 9.76 (8.69%) 2.91 (21.76%) 91.99 (112.64%) 51.89 (56.72%) 33.43 (90.59%) 54.73 (81.59%) 9.97 (8.84%)
0.15 48.45 (88.82%) 9.82 (9.35%) 2.92 (22.18%) 94.75 (119.02%) 51.88 (56.69%) 34.34 (95.78%) 55.69 (84.77%) 10.03 (9.50%)
0.20 49.47 (92.79%) 9.87 (9.91%) 2.95 (23.43%) 95.91 (121.71%) 52.79 (59.44%) 34.34 (95.78%) 56.76 (88.32%) 10.07 (9.93%)
0.25 49.39 (92.48%) 9.95 (10.80%) 2.94 (23.01%) 95.99 (121.89%) 52.98 (60.01%) 34.84 (98.63%) 56.79 (88.42%) 10.14 (10.70%)
SMPLer-X-B Clean 28.95 9.72 2.60 50.75 38.02 19.82 33.43 9.91
0.05 53.38 (84.39%) 10.61 (9.16%) 3.23 (24.23%) 106.97 (110.78%) 57.10 (50.18%) 36.21 (82.69%) 62.12 (85.82%) 10.82 (9.18%)
0.10 53.74 (85.63%) 10.66 (9.67%) 3.26 (25.38%) 106.27 (109.40%) 56.53 (48.68%) 35.96 (81.43%) 62.68 (87.50%) 10.87 (9.69%)
0.15 54.42 (87.98%) 10.69 (9.98%) 3.25 (25.00%) 110.59 (117.91%) 57.06 (50.08%) 36.81 (85.72%) 64.13 (91.83%) 10.85 (9.49%)
0.20 54.86 (89.50%) 10.80 (11.11%) 3.29 (26.54%) 110.25 (117.24%) 57.42 (51.03%) 36.46 (83.96%) 65.09 (94.71%) 10.98 (10.80%)
0.25 54.83 (89.40%) 10.75 (10.60%) 3.28 (26.15%) 111.13 (118.98%) 56.91 (49.68%) 36.14 (82.34%) 64.77 (93.75%) 10.94 (10.39%)
SMPLer-X-S Clean 32.28 10.05 2.83 57.23 42.80 22.26 37.38 10.27
0.05 50.16 (55.39%) 10.53 (4.78%) 3.33 (17.67%) 98.24 (71.66%) 59.06 (37.99%) 34.70 (55.88%) 58.62 (56.82%) 10.78 (4.97%)
0.10 50.28 (55.76%) 10.60 (5.47%) 3.32 (17.31%) 98.31 (71.78%) 59.02 (37.90%) 35.40 (59.03%) 58.52 (56.55%) 10.82 (5.36%)
0.15 50.69 (57.03%) 10.60 (5.47%) 3.35 (18.37%) 100.39 (75.41%) 59.54 (39.11%) 35.84 (61.01%) 58.99 (57.81%) 10.82 (5.36%)
0.20 51.24 (58.74%) 10.72 (6.67%) 3.38 (19.43%) 102.18 (78.54%) 59.70 (39.49%) 37.11 (66.71%) 60.07 (60.70%) 10.93 (6.43%)
0.25 50.90 (57.68%) 10.72 (6.67%) 3.40 (20.14%) 103.03 (80.03%) 59.98 (40.14%) 36.88 (65.68%) 59.87 (60.17%) 10.92 (6.33%)
Table 5: Performance comparison of different perturbation magnitude $\eta$ on state-of-the-art EHPS models on the PW3D dataset. The error growth rates are marked in gray. The maximum error and maximum error growth rate on each setting are highlighted underlined.
Model $\eta$ MPJPE (Body) $\downarrow$ (mm) PA MPJPE (Body) $\downarrow$ (mm)
SMPLer-X-H Clean 75.01 50.57
0.05 103.45 (37.91%) 65.40 (29.33%)
0.10 102.85 (37.12%) 65.41 (29.35%)
0.15 101.91 (35.86%) 65.20 (28.93%)
0.20 101.48 (35.29%) 65.10 (28.73%)
0.25 101.12 (34.81%) 64.93 (28.40%)
SMPLer-X-L Clean 75.85 50.67
0.05 104.53 (37.81%) 64.94 (28.16%)
0.10 104.81 (38.18%) 65.28 (28.83%)
0.15 104.64 (37.96%) 65.34 (28.95%)
0.20 104.55 (37.84%) 65.39 (29.05%)
0.25 104.17 (37.34%) 65.38 (29.03%)
SMPLer-X-B Clean 79.46 52.62
0.05 117.97 (48.46%) 73.94 (40.52%)
0.10 117.58 (47.97%) 73.88 (40.40%)
0.15 117.53 (47.91%) 73.59 (39.85%)
0.20 117.42 (47.77%) 73.43 (39.55%)
0.25 117.60 (48.00%) 73.36 (39.41%)
SMPLer-X-S Clean 82.67 56.65
0.05 117.78 (42.47%) 74.74 (31.93%)
0.10 118.07 (42.82%) 74.47 (31.46%)
0.15 118.16 (42.93%) 74.27 (31.10%)
0.20 118.13 (42.89%) 74.09 (30.79%)
0.25 118.32 (43.12%) 74.14 (30.87%)
Table 6: Performance comparison of different query times $t$ on state-of-the-art EHPS models on the PW3D dataset. The error growth rates are marked in gray. The maximum error and maximum error growth rate on each setting are highlighted underlined.
Model Attack MPJPE (Body) $\downarrow$ (mm) PA MPJPE (Body) $\downarrow$ (mm)
SMPLer-X-H Clean 75.01 50.57
1 82.28 (9.69%) 54.03 (6.84%)
3 103.45 (37.91%) 65.40 (29.33%)
5 127.04 (69.36%) 78.41 (55.05%)
7 153.67 (104.87%) 91.81 (81.55%)
9 177.11 (136.12%) 103.29 (104.25%)
SMPLer-X-L Clean 75.85 50.67
1 82.97 (9.39%) 54.29 (7.14%)
3 104.53 (37.81%) 64.94 (28.16%)
5 124.51 (64.15%) 76.01 (50.01%)
7 146.68 (93.38%) 87.63 (72.94%)
9 166.96 (120.12%) 97.83 (93.07%)
SMPLer-X-B Clean 79.46 52.62
1 87.34 (9.92%) 57.28 (8.86%)
3 117.97 (48.46%) 73.94 (40.52%)
5 140.52 (76.84%) 86.45 (64.29%)
7 157.20 (97.84%) 96.19 (82.80%)
9 171.32 (115.61%) 104.14 (97.91%)
SMPLer-X-S Clean 82.67 56.65
1 91.62 (10.83%) 61.12 (7.89%)
3 117.78 (42.47%) 74.74 (31.93%)
5 131.04 (58.51%) 80.92 (42.84%)
7 139.82 (69.13%) 85.24 (50.47%)
9 146.17 (76.81%) 88.47 (56.17%)
Table 7: Performance comparison of different query times $t$ on state-of-the-art EHPS models on the UBody dataset. The error growth rates are marked in gray. The maximum error and maximum error growth rate on each setting are highlighted underlined.
Model $t$ PA MPVPE $\downarrow$ (mm) MPVPE $\downarrow$ (mm) PA MPJPE $\downarrow$ (mm)
All Hands Face All Hands Face Body Hands
SMPLer-X-H Clean 24.64 8.47 2.24 41.22 30.88 16.62 29.29 8.64
1 30.74 (24.76%) 8.99 (6.14%) 2.41 (7.59%) 58.90 (42.89%) 36.80 (19.17%) 20.00 (20.34%) 35.67 (21.78%) 9.19 (6.37%)
3 43.94 (78.33%) 9.90 (16.88%) 2.80 (25.00%) 88.60 (114.94%) 48.69 (57.69%) 28.38 (70.76%) 50.48 (72.35%) 10.08 (16.67%)
5 55.14 (123.78%) 10.27 (21.25%) 3.22 (43.75%) 115.86 (181.08%) 57.41 (85.91%) 39.33 (136.64%) 63.62 (117.21%) 10.45 (20.95%)
7 64.32 (161.04%) 10.69 (26.21%) 3.59 (60.27%) 144.96 (251.67%) 66.02 (113.80%) 51.29 (208.60%) 76.34 (160.64%) 10.89 (26.04%)
9 70.83 (187.46%) 11.01 (29.99%) 3.89 (73.66%) 161.56 (291.95%) 73.72 (138.73%) 59.29 (256.74%) 83.76 (185.97%) 11.21 (29.75%)
SMPLer-X-L Clean 25.66 8.98 2.39 43.26 33.11 17.54 30.14 9.16
1 32.52 (26.73%) 9.05 (0.78%) 2.49 (4.18%) 59.09 (36.59%) 39.72 (19.96%) 21.18 (20.75%) 37.36 (23.95%) 9.23 (0.76%)
3 47.86 (86.52%) 9.83 (9.47%) 2.92 (22.18%) 91.46 (114.42%) 52.17 (57.57%) 33.01 (88.20%) 54.50 (80.82%) 10.03 (9.50%)
5 59.03 (130.05%) 10.34 (15.14%) 3.31 (38.49%) 120.70 (179.01%) 61.16 (84.72%) 44.91 (156.04%) 68.73 (128.04%) 10.53 (14.96%)
7 65.75 (156.24%) 10.62 (18.26%) 3.71 (55.23%) 146.06 (237.63%) 69.93 (111.21%) 55.43 (216.02%) 78.09 (159.09%) 10.82 (18.12%)
9 71.06 (176.93%) 11.01 (22.61%) 3.91 (63.60%) 157.47 (264.01%) 75.02 (126.58%) 60.80 (246.64%) 82.93 (175.15%) 11.23 (22.60%)
SMPLer-X-B Clean 28.95 9.72 2.60 50.75 38.02 19.82 33.43 9.91
1 36.04 (24.49%) 9.87 (1.54%) 2.83 (8.85%) 68.59 (35.15%) 43.18 (13.57%) 23.92 (20.69%) 41.09 (22.91%) 10.08 (1.72%)
3 53.38 (84.39%) 10.61 (9.16%) 3.23 (24.23%) 106.97 (110.78%) 57.10 (50.18%) 36.21 (82.69%) 62.12 (85.82%) 10.82 (9.18%)
5 64.52 (122.87%) 11.12 (14.40%) 3.60 (38.46%) 135.60 (167.19%) 68.92 (81.27%) 48.92 (146.82%) 76.61 (129.17%) 11.33 (14.33%)
7 71.25 (146.11%) 11.56 (18.93%) 3.96 (52.31%) 161.01 (217.26%) 79.57 (109.28%) 61.61 (210.85%) 84.00 (151.27%) 11.77 (18.77%)
9 74.96 (158.93%) 11.69 (20.27%) 4.25 (63.46%) 176.13 (247.05%) 85.22 (124.15%) 70.45 (255.45%) 88.87 (165.84%) 11.90 (20.08%)
SMPLer-X-S Clean 32.28 10.05 2.83 57.23 42.80 22.26 37.38 10.27
1 38.71 (19.92%) 10.11 (0.60%) 3.03 (7.07%) 70.78 (23.68%) 48.63 (13.62%) 25.55 (14.78%) 44.76 (19.74%) 10.34 (0.68%)
3 50.16 (55.39%) 10.53 (4.78%) 3.33 (17.67%) 98.24 (71.66%) 59.06 (37.99%) 34.70 (55.88%) 58.62 (56.82%) 10.78 (4.97%)
5 54.82 (69.83%) 10.82 (7.66%) 3.55 (25.44%) 114.22 (99.58%) 65.40 (52.80%) 44.03 (97.80%) 63.69 (70.39%) 11.06 (7.69%)
7 57.40 (77.82%) 10.96 (9.05%) 3.70 (30.74%) 125.21 (118.78%) 70.09 (63.76%) 51.48 (131.27%) 66.17 (77.02%) 11.20 (9.06%)
9 58.03 (79.77%) 11.01 (9.55%) 3.75 (32.51%) 130.08 (127.29%) 71.89 (67.97%) 54.82 (146.27%) 66.76 (78.60%) 11.25 (9.54%)
Effects of Different Perturbation Magnitude $\eta$. To gain a clearer understanding of the impact of varying perturbation magnitudes $\eta$ on the EHPS model, we conduct experiments using the SMPLer-X model and report the task-specific results for all tasks on the UBody and PW3D datasets, respectively. As illustrated in Tables. 4 and 5, different perturbation magnitudes $\eta$ result in comparable attack effectiveness. This phenomenon may be attributed to the model’s sensitivity to latent-space perturbations being concentrated along specific directions rather than being strictly dependent on the perturbation magnitude. As a result, increasing the magnitude does not substantially improve attack performance.
Effects of Different Query Times $t$. Similarly, we conduct an ablation study on the number of queries made to the EHPS model API, denoted as $t$, to examine its impact on the performance of adversarial attacks. We also report the task-specific results for all tasks on the UBody and PW3D datasets, respectively. As shown in Tables 6 and 7, the number of queries $t$ is positively correlated with the mean error growth rate, suggesting that increasing the number of queries enhances attack effectiveness. Notably, the error growth rate exceeds 100% on both MPVPE and PA MPJPE multiple estimation metrics, and even reaches a maximum of 255.45%. However, an excessive number of queries raises the attacker’s cost and may alert the platform or API owner. Therefore, although such attacks may not yield optimal performance, conducting them within a strict query budget enables sustained adversarial behavior.
Table 8: Quantitative comparison on UBody. The best and the second best results are highlighted and underlined.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
FGSM 46.69 0.98 0.13 2.64
PGD 40.79 0.92 0.23 7.24
ACA 29.01 0.99 0.01 11.26
DiffAttack 40.63 0.92 0.19 7.43
TBA 13.83 0.62 0.54 62.04
UBA (Ours) 47.15 0.98 0.10 1.89
Figure 5: Visualizing various adversarial samples for digital human generation on PW3D.
Figure 6: Visualizing various adversarial samples for digital human generation on UBody.
We adopt four standard image quality evaluation metrics, e.g., PSNR (Peak Signal-to-Noise Ratio), SSIM (Structural Similarity Index) [41], LPIPS (Learned Perceptual Image Patch Similarity) [43], and FID (Fréchet Inception Distance) [10], to assess the imperceptibility of the adversarial examples generated by our method. As presented in Table 8, our method achieves the best results in PSNR and FID, and second-best performance in SSIM and LPIPS. Furthermore, we conduct additional qualitative experiments by visualizing the generated digital humans from both clean and adversarial samples generated by various attack methods across various scenarios to substantiate our findings. As shown in Figures. 5 and 6, UBA causes a marked degradation in human pose and shape estimation accuracy, leading to evident distortions in the synthesized digital humans, and in some instances, rendering the generated poses entirely implausible. These results clearly demonstrate the adverse impact of UBA on the ability of EHPS models to accurately capture human pose.
Appendix C Social Impact
Positive Social Impacts. Although our work centers on adversarial attacks, its primary objective is to uncover vulnerabilities in digital human generation models and raise community awareness of the associated security risks. We seek to alert developers to these issues and promote the development of robust and secure digital human synthesis systems.
Negative Social Impacts. The experimental results validate the proposed method’s strong applicability and attack effectiveness. However, its potential misuse by malicious actors to compromise security-sensitive applications raises serious concerns about the robustness of digital human generation technologies.
Appendix D Limitations
While this work proposes an effective and imperceptible black-box attack scheme with strong real-world applicability, it also has limitations. Specifically, our evaluation focuses exclusively on attack efficacy, without considering the potential influence of standard input pre-processing defenses. This omission implies a potentially unrealistic assumption—that the API provider or model owner does not apply any input transformations or defense techniques (e.g., image compression, smoothing filters, JPEG encoding, or feature denoising) before inference. However, such defenses are often integrated into secure machine learning pipelines and may significantly degrade or nullify the adversarial perturbations generated by our method. Therefore, future research should systematically examine the effect of the proposed attack under various pre-processing defenses and adaptive countermeasures to more accurately assess its practical threat in real-world deployments.