### v2

# LatentStealth: Unnoticeable and Efficient Adversarial Attacks on Expressive Human Pose and Shape Estimation

Abstract
Expressive human pose and shape estimation (EHPS) plays a central role in digital human generation, particularly in live-streaming applications. However, most existing EHPS models focus primarily on minimizing estimation errors, with limited attention on potential security vulnerabilities, such as generating inappropriate content, violent actions, or racially offensive gestures and expressions. Current adversarial attacks on EHPS models often generate visually conspicuous perturbations, limiting their practicality and ability to expose real-world security threats. To address this limitation, we propose an unnoticeable adversarial method, termed LatentStealth, specifically tailored for EHPS models. The key idea is to exploit the structured latent representations of natural images as the medium for crafting perturbations. Instead of injecting noise directly into the pixel space, our method projects inputs into the latent space, where adversarial patterns are generated and progressively refined along optimized directions. This latent-space manipulation enables the attack to maintain high imperceptibility while preserving its effectiveness. Furthermore, as the optimization process is guided by only a small number of model output queries, the framework achieves competitive attack performance with low computational overhead, making it both practical and efficient for real-world scenarios. Extensive experiments on the 3DPW and UBody datasets demonstrate the superiority of LatentStealth, revealing critical vulnerabilities in current systems. These findings highlight the urgent need to address and mitigate security risks in digital human generation technologies.
I Introduction
Expressive human pose and shape estimation (EHPS) from monocular images or videos is fundamental to digital human modeling and key applications such as live streaming, virtual reality, and interactive gaming [24, 28, 43, 39, 2, 12, 11]. In these scenarios, EHPS systems play a critical role in animating realistic digital avatars capable of dynamically replicating human movements and facial expressions in real time, as shown in Fig. 1 (a). Although recent advancements have significantly enhanced estimation accuracy, the security vulnerabilities of such models remain largely underexplored. Growing evidence [36, 3, 31, 42] indicates that these systems are susceptible to adversarial perturbations. Although current attack methods have demonstrated promising effectiveness, their perturbations are typically visible and therefore easily detected, as shown in Fig. 1 (b), which has limited attention to this issue within the research community. However, if attacks are conducted in an imperceptible manner, they may undermine public safety. For instance, in digital human live streaming, such attacks could cause an avatar’s head to appear severed or clothing to fall off, leading to severe incidents involving horror or pornography. Therefore, raising awareness of this problem within both the research community and broader society is urgent.
Figure 1: Motivation and comparison of adversarial attacks on digital human generation. (a) A standard pipeline generates correct postures by disentangling pose, shape, and facial factors. (b) Existing attacks perturb the image space, causing visible artifacts and risky postures. (c) Our method manipulates the latent space, remaining visually imperceptible while still inducing erroneous posture estimation, revealing a more stealthy and realistic attack.
While these models exhibit strong performance, their robustness against adversarial attacks has received limited attention. From a methodological standpoint, existing EHPS studies primarily focus on reducing errors in body pose estimation [43, 39] and facial pose estimation [17, 26, 23, 2]. This is typically achieved using parametric human body models such as SMPL-X [28] or SMPLer-X [2], which leverage deep neural networks to capture the complex geometry and motion of the human body. TBA [22] was the first to propose a white-box transferable adversarial attack against EHPS. It utilizes the gradient of the SMPLer-X model to generate adversarial examples that significantly increase the estimation error across various EHPS models, thereby exposing potential security vulnerabilities. However, the perturbations produced by TBA are visually perceptible and fail to pass human inspection, which limits its applicability in real-world scenarios. Despite its effectiveness, such attacks have yet to attract sufficient attention from the research community. Considering these factors, we argue that an ideal adversarial attack on EHPS should meet three key criteria: i) the perturbation must be visually imperceptible; ii) the computational cost of the attacker should be minimal, aligning with realistic threat models; iii) the adversarial examples must achieve a high error growth rate. A substantial gap remains between current methods and these ideal standards.
To tackle this challenge, we propose a novel unnoticeable attack method, termed LatentStealth, against EHPS models. Inspired by [6, 30], LatentStealth perturbs the latent space of a pretrained Variational Autoencoder (VAE) [19, 10] by injecting low-magnitude Gaussian noise. Leveraging the nonlinear generative capabilities of the decoder of VAE [18, 34], our proposal can produce an initial adversarial example that is visually indistinguishable from the original input, as shown in Fig. 1 (c). The resulting pixel-wise difference is treated as an initial perturbation and subsequently refined through an iterative optimization process guided by a customized multi-task loss function. This loss simultaneously maximizes the discrepancy in the EHPS model’s predictions while minimizing perceptual deviation from the clean image. Moreover, since the attack operates under a strict test budget, its computational cost remains low, ensuring efficiency. We conduct extensive experiments on the 3DPW and UBody datasets to evaluate the effectiveness of LatentStealth. Results show that it achieves strong attack capability (increasing pose estimation errors by an average of 17.27%–58.21%), high stealthiness (achieving the best FID [9] and PSNR [38]), and low computational cost (requiring only three queries of the model outputs).
In conclusion, our main contributions are:
  • We propose the first unnoticeable adversarial attack method for EHPS systems under realistic constraints, which significantly degrades model performance while remaining visually undetectable. Moreover, our study highlights a critical yet underexplored security risk in digital human modeling.
  • We develop LatentStealth, a VAE-based attack that perturbs latent representations with low-magnitude noise and leverages the decoder’s nonlinear generative capacity to craft adversarial examples indistinguishable from clean inputs. This design fundamentally differs from pixel-space attacks by enhancing both stealthiness and naturalness.
  • We design a multi-task loss that jointly balances adversarial effectiveness and perceptual similarity. Coupled with an iterative refinement process constrained by a strict test budget, LatentStealth achieves strong attack performance with low computational overhead, ensuring practical applicability in real-world scenarios.
The remainder of this paper is organized as follows. Section II reviews the related work. Section III introduces the necessary background knowledge. Section IV details the proposed method. Section V reports comprehensive experimental results and comparisons with existing methods. Section VI concludes the paper. Section VII represents the limitation and future work. Section VIII discusses social impacts.
II Related Work
II-A Expressive Human Pose and Shape Estimation (EHPS)
EHPS is fundamental to digital human modeling, with applications in virtual environments, live broadcasting, avatars, and animation [37, 21, 29, 44, 27, 32]. Recent years have seen significant advances in improving the accuracy and realism of EHPS models. [28] introduced SMPL-X, a comprehensive 3D model incorporating articulated hands and expressive facial features for improved monocular pose capture. [26] proposed Hand4Whole, leveraging metacarpophalangeal joints to predict wrist rotations precisely while excluding full-body context for better finger estimation. [23] developed OSX, a one-stage framework using a component aware transformer to reconstruct whole-body meshes without post-processing. [2] presented SMPLer-X, a generalist foundation model trained on diverse datasets, achieving strong generalization and competitive performance across EHPS benchmarks. However, existing research has predominantly concentrated on enhancing estimation accuracy, with limited attention to the models’ vulnerability under attack threats.
II-B Adversarial Attack against Deep Neural Networks.
Adversarial attacks are commonly employed to assess the robustness of deep neural networks (DNNs) against malicious inputs [15, 8, 32, 13, 14, 41]. These attacks introduce subtle perturbations that mislead models while preserving the visual appearance of the input. [16] proposed TT3D, which integrates NeRF-based multi-view reconstruction with dual optimization of texture features and MLP parameters to generate natural-looking, transferable 3D adversarial examples. [5] presented the first content-aware, unrestricted adversarial attack based on diffusion models, enhancing transferability and realism by optimizing semantic perturbations in latent space. [4] proposed DiffAttack, a diffusion-based framework that achieves highly transferable and imperceptible attacks by jointly modulating cross-attention and self-attention mechanisms. Despite these advancements, the robustness of EHPS models remains largely underexplored. [22] proposed TBA consisting of a dual heterogeneous noise generator, the first EHPS-targeted attack; however, its visible perturbations limit its practicality. Consequently, TBA has failed to draw sufficient attention from the research community regarding EHPS security risks. To address this gap, we propose an unnoticeable attack method that reveals critical vulnerabilities in EHPS models.
Figure 2: Overview of the LatentStealth pipeline. The framework consists of two stages: a) noise injection in latent space, and b) noise enhancement.
III Preliminaries
The Main Pipeline of EHPS. Expressive human pose and shape estimation (EHPS) aims to reconstruct detailed 3D representations of the human body, face, and hands from monocular images. This capability underpins a wide array of applications, including virtual environments, live broadcasting, avatar generation, and animation. Formally, the ground-truth pose is denoted as $\alpha\in\mathbb{R}^{53\times 3}$, comprising the body pose $\alpha_{\text{body}}\in\mathbb{R}^{22\times 3}$, left- and right-hand poses $\alpha_{\text{lhand}},\alpha_{\text{rhand}}\in\mathbb{R}^{15\times 3}$, and the jaw pose $\alpha_{\text{jaw}}\in\mathbb{R}^{1\times 3}$. Additionally, individual-specific geometry and facial expressions are encoded by the shape vector $\beta\in\mathbb{R}^{10}$ and the expression vector $\gamma\in\mathbb{R}^{10}$, respectively. Let $\mathcal{P}$ denote the EHPS model; given a unseen monocular image $x$, the objective is to estimate the complete set of pose, shape, and expression parameters $(\alpha,\beta,\gamma)$, e.g., $\mathcal{P}(x)=(\alpha,\beta,\gamma)$. The estimation process minimizes the parameter reconstruction loss:
$\min_{{\alpha,\beta,\gamma}}\left(\|\alpha-\alpha_{gt}\|_{2}+\|\beta-\beta_{gt}\|_{2}+\|\gamma-\gamma_{gt}\|_{2}\right).$ (1)
Finally, 3D joints are rendered using the SMPL-X function $\mathcal{S}$ and a joint regressor $\mathcal{R}$, denoted as $\mathcal{R}(\mathcal{S}(\alpha,\beta,\gamma))$. $(\alpha_{gt},\beta_{gt},\gamma_{gt})$ denotes the ground-truth human pose and shape parameters.
Threat Model. This paper investigates adversarial attack techniques targeting EHPS models. Specifically, the attacker lacks access to any details of the model’s training process, including the dataset, architecture, and trianing trick, and is only permitted a limited number of acquring the model’s output. The generated adversarial examples are directly submitted to the model for inference. Other attack types with limited practical applicability (e.g., training-controlled or model modification attacks) require additional attacker capabilities and are beyond the scope of this study.
IV Method
IV-A Problem Formulation
As noted in previous sections, existing EHPS models primarily aim to reduce human pose and shape estimation errors, with limited emphasis on their vulnerability to attacks. Although prior work [22] introduced the first adversarial attack against EHPS, exposing its inherent fragility, its real-world applicability remains constrained due to the visible nature of adversarial perturbations. This limitation hinders the community’s full recognition of the intrinsic risks associated with EHPS. To tackle the challenge, we propose an unnoticeable attack method, termed LatentStealth, designed for realistic scenarios where the attacker can only access the model’s output under a strict query budget.
Attacker’s Goals. The LatentStealth is designed with two primary objectives: 1) effectiveness and 2) unnoticeability. Specifically, the effectiveness demands that the adversarial examples significantly increase errors in human pose and shape estimation, and the computational cost remains low. The unnoticeability ensures that EHPS model owners are unable to identify the adversarial examples. The formalization is as follows:
$\mathcal{P}(\hat{x})=(\hat{\alpha},\hat{\beta},\hat{\gamma}),\quad\mathcal{P}(x)=(\alpha,\beta,\gamma),$ (2)
$\begin{split}\max_{\hat{\alpha},\hat{\beta},\hat{\gamma}}\;&\Bigl(\|\hat{\alpha}-\alpha\|_{2}^{2}+\|\hat{\beta}-\beta\|_{2}^{2}+\|\hat{\gamma}-\gamma\|_{2}^{2}\Bigr)\\ \text{s.t.}\;&\|\hat{x}-x\|_{2}^{2}\leq\varepsilon,\end{split}$ (3)
where $x$ denotes a unseen monocular image (clean input image), $\hat{x}$ represents the corresponding adversarial sample image, and $\varepsilon$ indicates the perturbation threshold.
IV-B LatentStealth
Existing EHPS attacks often produce visible perturbations, limiting practicality. To overcome this, we propose LatentStealth, an unnoticeable attack method (as shown in Fig. 2) that perturbs the latent space of a pretrained Variational Autoencoder (VAE) [19] with low-magnitude noise and refines it via a multi-task loss under a strict test budget. LatentStealth achieves effective, efficient, and visually stealthy attacks, exposing critical yet overlooked security vulnerabilities in EHPS systems.
Noise Injection in Latent Space. Inspired by [33], who questioned the necessity of norm-bounded pixel constraints, we introduce adversarial perturbations in the latent space, thereby more naturally preserving the structural features of the input and improving the imperceptibility of the perturbations. Unlike conventional methods that manipulate input images directly [7, 25, 5, 4], this strategy targets internal representations that more directly govern model predictions. Importantly, perturbing the latent space allows for adversarial behavior while preserving high perceptual fidelity in the image space. To reduce attack cost and improve real-world applicability, we employ a publicly available pretrained VAE [19] without additional training or fine-tuning.
The VAE is trained by maximizing the evidence lower bound and utilizes the reparameterization trick to enable differentiable sampling as follows:
[…]
$z=\mu_{\phi}(x)+\sigma_{\phi}(x)\odot\epsilon_{1},\quad\epsilon_{1}\sim\mathcal{N}(0,I),$ (5)
where $\phi$ and $\theta$ are the encoder and decoder parameters, the encoder outputs $\mu_{\phi}(x)$ and $\sigma_{\phi}(x)$ as mean and standard deviation, $z\sim\mathcal{N}(0,I)$ is the standard Gaussian prior, and the decoder $\mathcal{G}_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{n}$ maps the $d$-dimensional latent space to the $n$-dimensional pixel space, yielding a smooth latent representation [20].
[…]
$\hat{x}\approx\mathcal{G}_{\theta}(z)+J_{\mathcal{G}_{\theta}}(z)\cdot(\eta\cdot\epsilon_{2}).$ (7)
Consequently, the perturbation $\Delta x=\hat{x}-x$ in the pixel space can be expressed as:
$\Delta x\approx J_{\mathcal{G}_{\theta}}(z)\cdot(\eta\cdot\epsilon_{2}).$ (8)
Specifically, the perturbation $\Delta x$ in the pixel space exhibits a linear relationship with the perturbation $\eta\cdot\epsilon_{2}$ in the latent space. Since the pre-trained model $\mathcal{G}_{\theta}$ is typically continuously differentiable, the magnitude of the perturbation is consequently constrained as follows:
[…]
If the perturbation magnitude $\eta$ is sufficiently small, the corresponding distortion in pixel space remains imperceptible, even in directions associated with large Jacobian values.
Similarly, a first-order approximation is applied to the target model:
$\begin{split}\mathcal{P}(\hat{x})&=\mathcal{P}\!\big(x+J_{\mathcal{G}_{\theta}}(z)\cdot(\eta\cdot\epsilon_{2})\big)\\ &\approx\mathcal{P}(x)+J_{\mathcal{P}}(x)\cdot J_{\mathcal{G}_{\theta}}(z)\cdot(\eta\cdot\epsilon_{2}).\end{split}$ (10)
We define the threshold for significant changes in the estimation parameters of the EHPS model as $\delta$. If the deviation exceeds this threshold, the model is considered to be misled, i.e., $\mathcal{P}(\hat{x})-\mathcal{P}(x)||_{2}\geq\delta$. According to Eq. 10, this leads to the inequality $J_{\mathcal{P}}(x)\cdot J_{\mathcal{G}_{\theta}}(z)\cdot(\eta\cdot\epsilon_{2})\geq\delta$.
By combining the Jacobian matrices into a single term $\mathbb{J}=J_{\mathcal{P}}(x)\cdot J_{\mathcal{G}_{\theta}}(z)$, the inequality can be reformulated as $\mathbb{J}\cdot(\eta\cdot\epsilon_{2})||_{2}\geq\delta$. Given that $\epsilon_{2}\sim\mathcal{N}(0,I)$, the expected perturbation is:
[…]
To minimize the perturbation magnitude $\eta$ while effectively crossing the decision boundary, the perturbation should be aligned with the right singular vector corresponding to the largest singular value:
$\mathbb{J}=U\Sigma V^{\top},\quad\Sigma=\text{diag}(\sigma_{1},\sigma_{2},\ldots),\quad\sigma_{1}\geq\sigma_{2}\geq\cdots$ (12)
The most sensitive direction corresponds to the first column vector $v_{1}$ of $V$. If $\epsilon_{2}$ is set to $\upsilon_{1}$, the condition for the minimum perturbation magnitude becomes:
$\mathbb{J}\cdot(\eta\cdot\upsilon_{1})||_{2}=\eta\cdot\sigma_{1}\geq\delta\Rightarrow\eta\geq\frac{\delta}{\sigma_{1}}.$ (13)
In summary, an attacker can amplify variations in sensitive features by injecting small perturbations into the latent space, causing the prediction to deviate rapidly from its original stable state and leading to significant errors.
Noise Enhancement. Inspired by Projected Gradient Descent (PGD) [25], we adopt an iterative refinement strategy to improve the effectiveness of adversarial examples, guided by feedback from the EHPS model. However, updating the noise in the latent space entails significant computational overhead. Specifically, the latent perturbation is defined as:
$\begin{split}\mathcal{P}(\mathcal{G}_{\theta}(z+\Delta{\hat{z}}^{(t)}))&=(\hat{\alpha}^{(t)},\hat{\beta}^{(t)},\hat{\gamma}^{(t)}),\\ \mathcal{P}(x)&=(\alpha,\beta,\gamma),\end{split}$ (14)
[…]
$\nabla_{\Delta{\hat{z}}^{(t)}}\mathcal{L}((\hat{\alpha}^{(t)},\hat{\beta}^{(t)},\hat{\gamma}^{(t)}),(\alpha,\beta,\gamma))=\frac{\partial\mathcal{L}}{\partial\mathcal{G}_{\theta}}\frac{\partial\mathcal{G}_{\theta}}{\partial\Delta{\hat{z}}^{(t)}},$ (16)
where $t$ denotes the iteration step, $\Delta{\hat{z}}^{(t)}$ represent the latent perturbation and its update at step $t$, $z$ is the latent representation of clean image, $(\alpha,\beta,\gamma)$ are the outputs obtained by feeding the clean image into $\mathcal{P}$, and $\xi_{z}$ is the learning rate for the gradient update in the latent space. The term $\frac{\partial\mathcal{G}_{\theta}}{\partial\Delta{\hat{z}}^{(t)}}$ denotes the Jacobian matrix of the decoder, whose computation is prohibitively expensive due to its high dimensionality ($n\times d$, often very large). Furthermore, each update necessitates a complete forward pass through the decoder network, resulting in substantial memory and computational consumption. Consequently, performing iterative updates in the latent space is highly inefficient (see Table V).
TABLE I: Performance comparison of different adversarial attack methods on state-of-the-art EHPS models on the UBody dataset. The error growth rates are marked in gray. The maximum error and maximum error growth rate on each setting are highlighted underlined.
Model Attack PA MPVPE $\downarrow$ ($m$) MPVPE $\downarrow$ ($m$) PA MPJPE $\downarrow$ ($m$)
All Hands Face All Hands Face Body Hands
SMPLer-X-H Clean 24.64 8.47 2.24 41.22 30.88 16.62 29.29 8.64
FGSM 29.83 (21.06%) 9.03 (6.61%) 2.41 (7.59%) 57.76 (40.13%) 36.03 (16.68%) 20.11 (21.00%) 34.82 (18.88%) 9.18 (6.25%)
PGD 29.59 (20.09%) 9.17 (8.26%) 2.49 (11.16%) 60.06 (45.71%) 34.50 (11.72%) 22.00 (32.37%) 34.55 (17.96%) 9.33 (7.99%)
ACA 26.84 (8.93%) 9.14 (7.91%) 2.40 (7.14%) 46.29 (12.30%) 34.12 (10.49%) 18.04 (8.54%) 31.39 (7.17%) 9.32 (7.87%)
DiffAttack 28.54 (15.83%) 9.32 (10.04%) 2.43 (8.48%) 53.41 (29.57%) 35.41 (14.67%) 19.70 (18.53%) 33.80 (15.40%) 9.48 (9.72%)
TBA 30.00 (21.75%) 9.09 (7.32%) 2.43 (8.48%) 58.13 (41.02%) 36.29 (17.52%) 21.19 (27.50%) 35.36 (20.72%) 9.25 (7.06%)
LatentStealth (Ours) 43.94 (78.33%) 9.90 (16.88%) 2.80 (25.00%) 88.60 (114.94%) 48.69 (57.69%) 28.38 (70.76%) 50.48 (72.35%) 10.08 (16.67%)
SMPLer-X-L Clean 25.66 8.98 2.39 43.26 33.11 17.54 30.14 9.16
FGSM 32.88 (28.14%) 9.32 (3.79%) 2.51 (5.02%) 63.99 (47.92%) 39.44 (19.12%) 22.52 (28.39%) 38.58 (28.00%) 9.49 (3.60%)
PGD 30.82 (20.11%) 9.36 (4.23%) 2.62 (9.62%) 63.07 (45.79%) 36.61 (10.57%) 22.84 (30.22%) 36.40 (20.77%) 9.52 (3.93%)
ACA 28.13 (9.63%) 9.69 (7.91%) 2.52 (5.44%) 49.39 (14.17%) 36.47 (10.15%) 19.18 (9.35%) 32.50 (7.83%) 9.86 (7.64%)
DiffAttack 30.65 (19.45%) 9.74 (8.46%) 2.49 (4.18%) 57.57 (33.08%) 37.96 (14.65%) 20.53 (17.05%) 36.32 (20.50%) 9.90 (8.08%)
TBA 33.50 (30.55%) 9.43 (5.01%) 2.69 (12.55%) 66.16 (52.94%) 40.50 (22.32%) 22.05 (25.71%) 39.01 (29.73%) 9.60 (4.80%)
LatentStealth (Ours) 47.86 (86.52%) 9.83 (9.47%) 2.92 (22.18%) 91.46 (114.42%) 52.17 (57.57%) 33.01 (88.20%) 54.50 (80.82%) 10.03 (9.50%)
SMPLer-X-B Clean 28.95 9.72 2.60 50.75 38.02 19.82 33.43 9.91
FGSM 38.06 (31.47%) 10.06 (3.50%) 2.83 (8.85%) 78.25 (54.19%) 45.42 (19.46%) 25.47 (28.51%) 43.83 (31.11%) 10.24 (3.33%)
PGD 34.56 (19.38%) 9.85 (1.34%) 2.87 (10.38%) 71.73 (41.34%) 41.46 (9.05%) 24.47 (23.46%) 40.16 (20.13%) 10.03 (1.21%)
ACA 31.48 (8.74%) 10.11 (4.01%) 2.74 (5.38%) 57.02 (12.35%) 41.31 (8.65%) 21.99 (10.95%) 36.16 (8.17%) 10.30 (3.94%)
DiffAttack 32.04 (10.67%) 10.12 (4.12%) 2.72 (4.62%) 58.23 (14.74%) 41.48 (9.10%) 21.79 (9.94%) 37.18 (11.22%) 10.30 (3.94%)
TBA 37.44 (29.33%) 9.80 (0.82%) 2.90 (11.54%) 76.41 (50.56%) 46.25 (21.65%) 25.63 (29.31%) 42.69 (27.70%) 10.04 (1.31%)
LatentStealth (Ours) 53.38 (84.39%) 10.61 (9.16%) 3.23 (24.23%) 106.97 (110.78%) 57.10 (50.18%) 36.21 (82.69%) 62.12 (85.82%) 10.82 (9.18%)
SMPLer-X-S Clean 32.28 10.05 2.83 57.23 42.80 22.26 37.38 10.27
FGSM 40.02 (23.98%) 10.73 (6.77%) 3.01 (6.36%) 82.44 (44.05%) 49.33 (15.26%) 29.30 (31.63%) 46.44 (24.24%) 10.96 (6.72%)
PGD 36.90 (14.31%) 10.63 (5.77%) 3.07 (8.48%) 78.03 (36.43%) 45.79 (6.99%) 26.85 (20.62%) 43.37 (16.02%) 10.89 (6.04%)
ACA 34.48 (6.82%) 10.53 (4.78%) 2.93 (3.53%) 61.99 (8.32%) 45.04 (5.23%) 23.63 (6.15%) 39.66 (6.10%) 10.72 (4.38%)
DiffAttack 34.27 (6.16%) 10.54 (4.88%) 2.89 (2.12%) 61.31 (7.13%) 45.25 (5.72%) 23.47 (5.44%) 39.43 (5.48%) 10.74 (4.58%)
TBA 39.52 (22.43%) 10.96 (9.05%) 3.17 (12.01%) 81.48 (42.37%) 49.46 (15.56%) 28.72 (29.02%) 46.06 (23.22%) 11.25 (9.54%)
LatentStealth (Ours) 50.16 (55.39%) 10.53 (4.78%) 3.33 (17.67%) 98.24 (71.66%) 59.06 (37.99%) 34.70 (55.88%) 58.62 (56.82%) 10.78 (4.97%)
OSX Clean 40.85 9.37 3.54 86.36 42.72 25.72 49.26 9.62
FGSM 45.68 (11.82%) 9.86 (5.23%) 3.29 (-7.06%) 104.51 (21.02%) 46.71 (9.34%) 33.64 (30.84%) 54.17 (9.97%) 10.13 (5.30%)
PGD 42.47 (3.97%) 9.68 (3.31%) 3.00 (-15.25%) 97.86 (13.32%) 43.89 (2.74%) 28.78 (11.94%) 51.11 (3.76%) 9.95 (3.43%)
ACA 41.93 (2.64%) 9.66 (3.09%) 3.48 (-1.69%) 89.10 (3.17%) 45.02 (5.38%) 27.05 (5.21%) 49.92 (1.34%) 9.94 (3.33%)
DiffAttack 40.66 (-0.47%) 9.67 (3.20%) 3.11 (-12.15%) 88.86 (2.89%) 43.99 (2.97%) 26.04 (1.28%) 48.64 (-1.26%) 9.95 (3.43%)
TBA 41.37 (1.27%) 9.67 (3.20%) 2.98 (-15.82%) 97.25 (12.61%) 45.37 (6.20%) 28.05 (9.10%) 49.57 (0.63%) 9.94 (3.33%)
LatentStealth (Ours) 60.36 (47.76%) 10.44 (11.42%) 3.86 (9.04%) 134.39 (55.62%) 55.82 (30.66%) 47.73 (85.65%) 75.17 (52.60%) 10.73 (11.54%)
Hand4Whole Clean 41.04 7.85 2.96 93.81 35.60 30.30 49.80 7.97
FGSM 41.03 (-0.02%) 7.85 (0.00%) 2.96 (0.00%) 93.82 (0.01%) 35.61 (0.03%) 30.30 (0.00%) 49.80 (0.00%) 7.97 (0.00%)
PGD 41.06 (0.05%) 7.85 (0.00%) 2.96 (0.00%) 93.87 (0.06%) 35.62 (0.06%) 30.31 (0.03%) 49.82 (0.04%) 7.97 (0.00%)
ACA 41.40 (0.88%) 7.83 (-0.25%) 2.98 (0.68%) 95.15 (1.43%) 36.02 (1.18%) 30.71 (1.35%) 50.34 (1.08%) 7.95 (-0.25%)
DiffAttack 41.10 (0.15%) 7.99(1.78%) 2.91 (-1.69%) 94.79 (1.04%) 36.02 (1.18%) 35.11 (15.87%) 50.12 (0.64%) 8.10 (1.63%)
TBA 43.72 (6.53%) 8.10 (3.18%) 3.00 (1.35%) 104.31 (11.19%) 36.57 (2.72%) 34.52 (13.93%) 53.22 (6.87%) 8.22 (3.14%)
LatentStealth (Ours) 56.59 (37.89%) 9.74 (24.08%) 3.52 (18.92%) 128.57 (37.05%) 54.02 (51.74%) 41.83 (38.05%) 71.67 (43.92%) 9.93 (24.59%)
To address the aforementioned issue, we propose enhancing the noise directly in the pixel space. The initial pixel-space perturbation is defined as:
$\Delta x^{(0)}=x^{(0)}-x,\quad x^{(0)}=\mathcal{G}_{\theta}(z+\eta\cdot\epsilon_{2}).$ (17)
We perform iterative updates to the perturbation in pixel space as follows:
$\mathcal{P}(x+\Delta x^{(t)})=(\hat{\alpha}^{(t)},\hat{\beta}^{(t)},\hat{\gamma}^{(t)}),$ (18)
$\begin{split}\Delta x^{(t+1)}&=\Delta x^{(t)}+\\ &\quad\xi_{\Delta x}\cdot\nabla_{\Delta x^{(t)}}\mathcal{L}((\hat{\alpha}^{(t)},\hat{\beta}^{(t)},\hat{\gamma}^{(t)}),(\alpha,\beta,\gamma)),\end{split}$ (19)
$\nabla_{\Delta x^{(t)}}\mathcal{L}((\hat{\alpha}^{(t)},\hat{\beta}^{(t)},\hat{\gamma}^{(t)}),(\alpha,\beta,\gamma))=\frac{\partial\mathcal{L}}{\partial(x+\Delta x^{(t)})},$ (20)
where $\xi_{\Delta x}$ denotes the learning rate for gradient updates in pixel space. Importantly, computing gradients in pixel space does not require backpropagation through the decoder network. Instead, perturbations are applied directly to the original image $x$, thereby eliminating the need to compute the high-dimensional Jacobian transformation from latent to pixel space. This approach substantially reduces both the computational complexity and memory usage of each gradient update step.
Furthermore, to ensure strong adversarial strength while maintaining the perturbation’s imperceptibility, we incorporate a regularization term into the optimization process. To this end, we introduce a novel multi-task loss function to jointly achieve these objectives:
$\mathcal{L}$ $=-\mathbb{E}_{\Delta x}\Big[\|\hat{\alpha}^{(t)}-\alpha\|_{2}^{2}+\|\hat{\beta}^{(t)}-\beta\|_{2}^{2}+\|\hat{\gamma}^{(t)}-\gamma\|_{2}^{2}\Big]$ (21)
$\quad+\frac{1}{w\times h}\sum_{i=0}^{w-1}\sum_{j=0}^{h-1}\big\|\Delta x^{(t)}(i,j)\big\|_{2}^{2}.$
where $w$ and $h$ denote the width and height of $x$, respectively. The first term encourages the perturbed pose parameters to diverge from those of $\mathcal{P}(x)$, thereby enhancing the effectiveness of the attack. The second term imposes a strict constraint on pixel-level variations to ensure the imperceptibility of the perturbation, effectively serving as a regularization term. According the Eqs 18, 20 and 21, the update rule for the pixel-space perturbation is formulated as:
$\Delta x^{(t+1)}=\Delta x^{(t)}+\xi_{\Delta x}\cdot(\frac{\partial\mathcal{L}}{\partial(x+\Delta x^{(t)})}-\lambda\cdot x^{(t)}),$ (22)
where $\lambda$ is a hyperparameter that balances attack strength and imperceptibility. It is evident that the computational complexity of updates in latent space is $\mathcal{O}(n\times d)$, whereas the complexity of pixel-space updates is only $\mathcal{O}(n)$. Therefore, the proposed noise enhancement strategy in pixel space incurs significantly lower computational overhead (see the Method C and LatentStealth in Table V). To provide a clearer understanding of the LatentStealth pipeline, we present a detailed the step of generating adversarial samples; see Algorithm 1.
Algorithm 1 Generation of Adversarial Sample
Input: Original Image $x$, Random Gaussian Noise $\epsilon_{1}$, $\epsilon_{2}$
Output: Adversarial Sample $\hat{x}$
Setting: Encoder of VAE $\mathcal{E}_{\phi}$, Decoder of VAE $\mathcal{G}_{\theta}$, EHPS model $\mathcal{P}$, Mean value $\mu_{\phi}$, Variance $\sigma_{\phi}^{2}$, Number of iterative queries $N$, Step size $\upsilon$, Noise magnitude $\eta$, Learning rate $\zeta$, Latent representation $z$, Multi-task loss function $\mathcal{L}$
1: Loading parameters $\phi$, $\theta$ of VAE
[…]
7:for $i$ = 1 to NN do
8:  $\hat{x}=x+\Delta x$
[…]
10:  $\mathcal{L}=\mathcal{L}(\hat{\alpha},\hat{\beta},\hat{\gamma},\hat{x},\alpha,\beta,\gamma,x)$
11:  Update
12:  $\Delta x=\Delta x-\zeta\nabla_{\Delta x}\mathcal{L}$
13:end for
14:return $\hat{x}=\Delta x+x$
V Experiments
V-A Experimental Settings
Datasets and EHPS Models. We evaluate our method on two benchmarks: 3DPW [35], with in-the-wild RGB videos and 3D SMPL pose annotations, and UBody [23], a large-scale upper-body mesh dataset with 1.05 million annotated images across 15 scenarios. For evaluation, we test LatentStealth on the three most representative EHPS models, including Hand4Whole [26], OSX [23], and SMPLer-X [2], which provides four ViT-based pre-trained variants, denoted as SMPLer-X-M, where M denotes the size of the Vision Transformer (ViT) backbone [1] (M $\in$ S, B, L, H).
TABLE II: Robustness evaluation of the state-of-the-art EHPS model SMPLer-X-H on the UBody dataset under adversarial perturbations from four types of random noise and a central white patch.
Attack PA MPVPE $\downarrow$ ($m$) MPVPE $\downarrow$ ($m$) PA MPJPE $\downarrow$ ($m$)
All Hands Face All Hands Face Body Hands
Clean 24.64 8.47 2.24 41.22 30.88 16.62 29.29 8.64
Random Noise $\uparrow$ 24.66 8.47 2.24 41.26 30.88 16.62 29.31 8.64
Random Noise $\uparrow\uparrow$ 24.86 8.44 2.24 41.98 31.06 16.76 29.57 8.61
Random Noise $\uparrow\uparrow\uparrow$ 25.34 8.43 2.25 44.29 31.44 17.07 30.22 8.59
Random Noise $\uparrow\uparrow\uparrow\uparrow$ 25.90 8.42 2.27 47.19 31.81 17.52 30.99 8.58
Patch 24.63 8.48 2.24 41.20 30.86 16.61 29.28 8.65
LatentStealth (Ours) 43.94 9.90 2.80 88.60 48.69 28.38 50.48 10.08
TABLE III: Performance comparison of different adversarial attack methods on state-of-the-art EHPS models on the 3DPW dataset. The error growth rates are marked in gray. The maximum error and maximum error growth rate on each setting are highlighted underlined.
Model Attack MPJPE (Body) $\downarrow$ ($m$) PA MPJPE (Body) $\downarrow$ ($m$)
SMPLer-X-H Clean 75.01 50.57
FGSM 84.40 (12.52%) 56.04 (10.82%)
PGD 88.70 (18.25%) 57.17 (13.05%)
ACA 77.30 (3.05%) 52.47 (3.76%)
DiffAttack 80.46 (7.27%) 54.26 (7.30%)
TBA 88.43 (17.89%) 59.39 (17.44%)
LatentStealth (Ours) 103.45 (37.91%) 65.40 (29.33%)
SMPLer-X-L Clean 75.85 50.67
FGSM 85.05 (12.13%) 56.20 (10.91%)
PGD 90.16 (18.86%) 57.30 (12.12%)
ACA 78.53 (3.53%) 53.02 (4.64%)
DiffAttack 78.70 (3.76%) 53.25 (5.09%)
TBA 91.63 (20.80%) 60.54 (19.48%)
LatentStealth (Ours) 104.53 (37.81%) 64.94 (28.16%)
SMPLer-X-B Clean 79.46 52.62
FGSM 94.04 (18.35%) 60.97 (15.87%)
PGD 94.45 (18.86%) 59.00 (12.12%)
ACA 82.68 (4.05%) 55.60 (5.66%)
DiffAttack 82.93 (4.37%) 56.31 (7.01%)
TBA 97.50 (22.70%) 63.55 (20.77%)
LatentStealth (Ours) 117.97 (48.46%) 73.94 (40.52%)
SMPLer-X-S Clean 82.67 56.65
FGSM 100.05 (21.02%) 65.32 (15.30%)
PGD 98.65 (19.33%) 63.19 (11.54%)
ACA 85.52 (3.45%) 58.86 (3.90%)
DiffAttack 85.61 (3.56%) 58.97 (4.10%)
TBA 108.15 (30.82%) 69.63 (22.91%)
LatentStealth (Ours) 117.78 (42.47%) 74.74 (31.93%)
OSX Clean 94.32 63.87
FGSM 102.09 (8.24%) 68.13 (6.67%)
PGD 99.58 (5.58%) 66.01 (3.35%)
ACA 98.05 (3.95%) 66.39 (3.95%)
DiffAttack 97.78 (3.67%) 66.10 (3.49%)
TBA 97.33 (3.19%) 65.32 (3.49%)
LatentStealth (Ours) 116.89 (23.93%) 75.34 (17.96%)
Hand4Whole Clean 100.66 67.93
FGSM 100.66 (0.00%) 67.93 (0.00%)
PGD 100.66 (0.00%) 67.93 (0.00%)
ACA 100.70 (0.04%) 67.95 (0.03%)
DiffAttack 100.54 (-0.12%) 67.68 (-0.37%)
TBA 115.07 (14.32%) 74.18 (9.20%)
LatentStealth (Ours) 128.72 (27.88%) 80.10 (17.92%)
Baseline and Evaluation Metrics. Currently, only one adversarial attack targets EHPS. To evaluate LatentStealth, we compare it with state-of-the-art methods in computer vision, including classical attacks (FGSM [7], PGD [25]), diffusion-based approaches (ACA [5], DiffAttack [4]), and the EHPS-specific TBA [22]. We report MPJPE and MPVPE for 3D joint and mesh accuracy, and PA-MPJPE/PA-MPVPE for alignment-aware evaluation, following prior EHPS benchmarks [26, 23, 2]. All metrics are in millimeters (mm).
Implementation Details. The experiments are conducted on a system running Ubuntu 20.04.6 LTS, using PyTorch framework, with a single NVIDIA L20 GPU. The VAE model employed is stabilityai/sd-vae-ft-mse from Hugging Face  [^1] An open source platform. https://huggingface.co/stabilityai/sd-vae-ft-mse. The maximum perturbation magnitude is set to $\eta=0.05$ with a step size of 0.01, and an $l_{\infty}$ constraint was enforced to ensure that changes in individual pixel values did not exceed 0.05. For baseline methods involving iterative optimization, the number of iterations is fixed at 20. For LatentStealth, the number of queries is set to $t=3$, and the balancing hyperparameter is fixed at $\lambda=0.5$.
V-B Robustness of EHPS Models
To evaluate the robustness of the EHPS model, we attack the state-of-the-art SMPLer-X-H using adversarial samples generated from four types of random noise and a central white patch on the UBody dataset. As shown in Table II, the performance metrics under these attacks deviate negligibly ($\leq$ 1%) from those on clean samples, indicating that ordinary noise does not meaningfully degrade EHPS model performance due to robust generalization from large, diverse training data. In contrast, our proposed LatentStealth successfully compromises SMPLer-X-H, revealing significant vulnerabilities to strategically crafted adversarial samples.
V-C Attack Performance on EHPS Models
Figure 3: Visualizing various adversarial samples for digital human generation on UBody.
TABLE IV: Quantitative comparison on UBody. The best result is highlighted.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ $_{2} \downarrow LL_{\infty}$ $\downarrow$
FGSM 26.91 0.45 0.48 52.01 34.72 0.10
PGD 24.90 0.33 0.48 48.96 43.72 0.10
ACA 26.99 0.92 0.21 18.18 35.48 0.11
DiffAttack 40.72 0.91 0.19 7.19 7.07 0.05
TBA 13.83 0.62 0.54 62.04 157.52 0.48
LatentStealth (Ours) 47.15 0.98 0.10 1.89 3.37 0.01
Figure 4: Visualizing various adversarial samples for digital human generation on PW3D.
We first evaluate the effectiveness of LatentStealth on the UBody dataset, which includes full-body, hand, and facial regions of digital humans. As shown in Table I, across all tasks, the adversarial examples generated by LatentStealth cause the most significant degradation in the accuracy of all EHPS models, compared to other attack methods. Specifically, the average increase in estimation error ranged from 25.13% to 58.21%. Significantly, for the SMPLer-X-H model, the MPVPE (All) (sixth column) increased from 41.22 mm to 88.60 mm, corresponding to a 114.94% error growth rate.
TABLE V: Ablation study of different adversarial attack frameworks on state-of-the-art EHPS models on the UBody dataset. The error growth rates are marked in gray. The maximum error and maximum error growth rate on each setting are highlighted underlined.
Model Method PA MPVPE $\downarrow$ ($m$) MPVPE $\downarrow$ ($m$) PA MPJPE $\downarrow$ ($m$) Memory
All Hands Face All Hands Face Body Hands
SMPLer-X-H Clean 24.64 8.47 2.24 41.22 30.88 16.62 29.29 8.64 -
A 32.41 (31.53%) 8.98 (6.02%) 2.41 (7.59%) 63.43 (53.88%) 38.65 (25.16%) 20.16 (21.30%) 38.29 (30.73%) 9.12 (5.56%) 9.93 GB
B 26.27 (6.62%) 8.58 (1.30%) 2.22 (-0.89%) 47.68 (15.67%) 32.28 (4.53%) 18.07 (8.72%) 31.27 (6.76%) 8.74 (1.16%) 7.00 GB
C Out of Memory $\geq$ 45 GB
LatentStealth (Ours) 43.94 (78.33%) 9.90 (16.88%) 2.80 (25.00%) 88.60 (114.94%) 48.69 (57.69%) 28.38 (70.76%) 50.48 (72.35%) 10.08 (16.67%) 12.83 GB
SMPLer-X-L Clean 25.66 8.98 2.39 43.26 33.11 17.54 30.14 9.16 -
A 35.64 (38.89%) 9.18 (2.23%) 2.48 (3.77%) 69.94 (61.97%) 41.61 (25.67%) 23.45 (33.69%) 42.31 (40.38%) 9.35 (2.07%) 6.20 GB
B 27.53 (7.29%) 8.99 (0.11%) 2.42 (1.26%) 50.45 (16.62%) 34.09 (2.96%) 19.09 (8.84%) 32.31 (7.20%) 9.16 (0.00%) 6.77 GB
C 27.24 (6.16%) 8.99 (0.11%) 2.40 (0.42%) 47.74 (10.36%) 34.21 (3.32%) 18.37 (4.73%) 32.05 (6.34%) 9.17 (0.11%) 43.79 GB
LatentStealth (Ours) 47.86 (86.52%) 9.83 (9.47%) 2.92 (22.18%) 91.46 (114.42%) 52.17 (57.57%) 33.01 (88.20%) 54.50 (80.82%) 10.03 (9.50%) 8.25 GB
SMPLer-X-B Clean 28.95 9.72 2.60 50.75 38.02 19.82 33.43 9.91 -
A 37.01 (27.84%) 9.68 (-0.41%) 2.71 (4.23%) 74.09 (45.99%) 44.58 (17.25%) 24.02 (21.19%) 43.11 (28.96%) 9.86 (-0.50%) 3.45 GB
B 31.48 (8.74%) 9.49 (-2.37%) 2.59 (-0.38%) 56.79 (11.90%) 39.01 (2.60%) 21.56 (8.78%) 37.00 (10.68%) 9.67 (-2.42%) 6.24 GB
C 27.24 (-5.91%) 8.99 (-7.51%) 2.40 (-7.69%) 47.74 (-5.93%) 34.21 (-10.02%) 18.37 (-7.32%) 32.05 (-4.13%) 9.17 (-7.47%) 43.63 GB
LatentStealth (Ours) 53.38 (84.39%) 10.61 (9.16%) 3.23 (24.23%) 106.97 (110.78%) 57.10 (50.18%) 36.21 (82.69%) 62.12 (85.82%) 10.82 (9.18%) 8.32 GB
SMPLer-X-S Clean 32.28 10.05 2.83 57.23 42.80 22.26 37.38 10.27 -
A 38.80 (20.20%) 10.38 (3.28%) 2.85 (0.71%) 76.07 (32.92%) 48.19 (12.59%) 25.81 (15.95%) 45.44 (21.56%) 10.60 (3.21%) 2.73 GB
B 35.74 (10.72%) 10.02 (-0.30%) 2.88 (1.77%) 67.46 (17.88%) 44.61 (4.23%) 23.70 (6.47%) 41.41 (10.78%) 10.26 (-0.10%) 7.04 GB
C 33.89 (4.99%) 10.10 (0.50%) 2.85 (0.71%) 60.28 (5.33%) 44.20 (3.27%) 22.74 (2.16%) 39.21 (4.90%) 10.34 (0.68%) 43.33 GB
LatentStealth (Ours) 50.16 (55.39%) 10.53 (4.78%) 3.33 (17.67%) 98.24 (71.66%) 59.06 (37.99%) 34.70 (55.88%) 58.62 (56.82%) 10.78 (4.97%) 7.47 GB
(a) UBody
(b) PW3D
Figure 5: Performance of different perturbation magnitude $\eta$.
(a) UBody
(b) PW3D
Figure 6: Performance of different test times $t$.
We further evaluate LatentStealth on the 3DPW dataset, focusing on how adversarial examples impact the estimation of body parts across different EHPS models. As shown in Table III, LatentStealth consistently causes the greatest reduction in model accuracy. Especially, LatentStealth yields a minimum error growth rate of 17.92% (increasing PA MPJPE from 67.93 mm to 80.10 mm on the Hand4Whole model) and a maximum of 48.46% (raising MPJPE from 79.46 mm to 118.00 mm on the SMPLer-X-B model), with average increases ranging from 17.27% to 44.49%. These results demonstrate that LatentStealth is highly destructive. Remarkably, its computational cost remains low, ensuring efficiency. Interestingly, we observe negative error growth rates in certain cases, such as PA MPVPE (Face) under the OSX model on UBody and several other metrics under specific settings. This finding suggests that while adversarial attacks generally impair EHPS performance, they may, in rare instances, also enhance estimation accuracy.
To further illustrate the effectiveness of LatentStealth, we conduct qualitative evaluations by visualizing the digital humans generated from both clean and adversarial samples produced by various attack methods. As shown in Figs. 3 and 4, LatentStealth leads to noticeable degradation in pose and shape estimation accuracy, resulting in significant deviations in the synthesized digital humans. These visual results clearly highlight the adverse impact of LatentStealth on the EHPS models’ ability to accurately reconstruct human poses.
V-D Stealthiness
We adopt six standard image quality evaluation metrics, e.g., PSNR (Peak Signal-to-Noise Ratio), SSIM (Structural Similarity Index) [38], LPIPS (Learned Perceptual Image Patch Similarity) [40], FID (Fréchet Inception Distance) [9], $L_{2}$ error, and $L_{\infty}$ error, to assess the imperceptibility of the adversarial examples generated by our method. As presented in Table IV, our method achieves the best results in all metrics. Furthermore, we conduct additional qualitative experiments by visualizing the generated digital humans from both clean and adversarial samples generated by various attack methods across various scenarios to substantiate our findings. As shown in Figs. 3 and 4, LatentStealth causes a marked degradation in human pose and shape estimation accuracy, leading to evident distortions in the synthesized digital humans, and in some instances, rendering the generated poses entirely implausible. These results clearly demonstrate the adverse impact of LatentStealth on the ability of EHPS models to accurately capture human pose.
V-E Ablation Study
Effects of Different Noise Injection Strategies. To further assess LatentStealth, we compare three alternative noise injection strategies within its framework: A) adding random noise in the pixel space and updating it iteratively via EHPS test; B) injecting random noise into the latent space after VAE encoding, then resampling and decoding to produce adversarial examples; C) injecting random noise into the VAE latent space, followed by resampling, decoding, and iterative updates via EHPS queries. We evaluate these strategies on SMPLer-X using the UBody dataset and report GPU memory usage. As shown in Table V, LatentStealth achieves strong attack performance with only 7.47–12.83 GB of memory, demonstrating its effectiveness and efficiency without requiring extensive resources.
Effects of Different Perturbation Magnitude $\eta$. To examine the effect of perturbation magnitude $\eta$ on EHPS models, we evaluate SMPLer-X and report mean error growth rates on the UBody and PW3D datasets. As shown in Fig. 5, different values of $\eta$ yield similar attack effectiveness. This indicates that the model’s sensitivity to latent-space perturbations is concentrated along specific directions, rather than determined solely by magnitude. Consequently, increasing $\eta$ does not substantially enhance attack performance.
Effects of Different Test Times $t$. Similarly, we conduct an ablation study on the number of tests $t$ to the EHPS model, to evaluate its effect on attack performance. Due to space limits, we report only mean error growth rates. As shown in Fig. 6, $t$ is positively correlated with error growth, indicating that more tests enhance attack effectiveness. When $t=9$, the mean error growth rates for SMPLer-X-B, SMPLer-X-L, and SMPLer-X-H all exceed 100%. However, excessive tests raise computational cost and reduce efficiency. Although LatentStealth may not perform optimally at $t=3$, a low-cost attack remains attractive to adversaries.
VI Conclusion
This paper proposes a novel unnoticeable attack method, termed LatentStealth. The proposed method leverages the structured and continuous properties of the latent space in VAE. Specifically, input samples are first projected into the latent space, perturbed with low-magnitude noise, and then decoded back to image space to generate adversarial variants that remain close to the original inputs. The perturbation pattern is iteratively optimized under strict query-budget constraints, with refinement guided by feedback from the EHPS model to ensure computational efficiency. To further balance attack effectiveness and visual imperceptibility, the multi-task loss is employed to jointly optimize adversarial success and perceptual similarity. Experiments show that LatentStealth achieves high attack success, strong imperceptibility, and low computational cost. Beyond EHPS models, this work highlights broader vulnerabilities in EHPS systems and raises awareness in the research community.
VII Limitation & Future Work
While this work proposes an effective and imperceptible attack scheme with strong real-world applicability, it also has limitations. Specifically, our evaluation focuses exclusively on attack efficacy, without considering the potential influence of standard input pre-processing defenses. This omission implies a potentially unrealistic assumption—that the API provider or model owner does not apply any input transformations or defense techniques (e.g., image compression, smoothing filters, JPEG encoding, or feature denoising) before inference. However, such defenses are often integrated into secure machine learning pipelines and may significantly degrade or nullify the adversarial perturbations generated by our method. Therefore, future research should systematically examine the effect of the proposed attack under various pre-processing defenses and adaptive countermeasures to more accurately assess its practical threat in real-world deployments.
VIII Social Impact
Positive Social Impacts. Although our work centers on adversarial attacks, its primary objective is to uncover vulnerabilities in digital human generation models and raise community awareness of the associated security risks. We seek to alert developers to these issues and promote the development of robust and secure digital human synthesis systems.
Negative Social Impacts. The experimental results validate the proposed method’s strong applicability and attack effectiveness. However, its potential misuse by malicious actors to compromise security-sensitive applications raises serious concerns about the robustness of digital human generation technologies. For instance, in digital human live streaming, such attacks could cause an avatar’s head to appear severed or clothing to fall off, leading to severe incidents involving horror or pornography.